### v1

# Asymptotic Behavior of Rupture Solutions for the Elliptic MEMS Equation with Hénon-Type Term

Abstract: For an elliptic MEMS equation with Hénon-type $\Delta u=\dfrac{\lambda|x|^{\alpha}}{u^{p}}+F$. we study rupture solutions (i.e. solutions for which $u(x_{0})=0$ at some point $x_{0}$, also $x_{0}$ is called a rupture point). In this paper we focus on the special case where the rupture occurs at the origin. According to the different Hénon-type exponents $\alpha$, we analyze the asymptotic behavior of such solutions near the origin, derive a full asymptotic expansion of arbitrary order in a neighborhood of the origin. Moreover, with respect to radial solution and non-radial solution with asymptotic radial condition, we prove both of them exist near the rupture point by constrcuting it.
Keywords:Asymptotic expansions, Rupture point, Hénon-type.
MSC:
1 Introduction
Micro-electro-mechanical systems (MEMS) are devices that integrate miniature components by combining electrostatic effects with microfabrication technology. Nowadays MEMS devices are widely used in electronic equipment, aerospace engineering and medical applications. With the rapid development of high-tech fields such as artificial intelligence and ultra-precision machining, the performance requirements on MEMS have been increasing; however, electrostatic MEMS typically suffer from pull-in instability.
Taking the electrostatically actuated elastic membrane system considered in this paper as a prototype, we proceed as follows. As illustrated in Figure 1, a deformable elastic membrane is put above a fixed ground plate. When a voltage is applied, the membrane is attracted towards the plate. As the voltage increases, the membrane eventually touches the plate. This so-called pull-in (or touchdown) instability may affect the stable operation of high-precision microelectronic devices and may even cause irreversible damage. On the other hand, in some applications such contact is actually desired. For instance, the triggering of air bag restraint system or the operation of printers. In this work we focus on the profile of the membrane near the touchdown point at the moment when contact occurs.
Figure 1:
We consider the following model. Let $u$ denote the distance between the membrane and the ground plate, $\lambda$ the applied voltage, and $|x|^{\alpha}$ the dielectric profile on the membrane (representing the spatial inhomogeneity induced by the variation of the dielectric constant). We are interested in the situation where the membrane touches the ground plate at $x=0$. In this case the governing main equation can be written as $\Delta u=\dfrac{\lambda|x|^{\alpha}}{u^{2}},x\in\mathbb{R}^{2}\setminus\{0\}$ ( $\lambda>0$ and $\alpha>-2$). In this paper we are interested in a more general equation. More precisely, we generalize the negative exponent $p=2$ in the right-hand side of the equation to a general $p>0$ with $p>1/2\alpha$, introduce an additional force term $F$, and extend the above two-dimensional model to the NN-dimensional setting($N\geq 1$). In the present work we focus on the following problem:
$\begin{cases}\Delta u=\dfrac{\lambda|x|^{\alpha}}{u^{p}}+F,&x\in\mathbb{R}^{N}\setminus\{0\},\\[4.0pt] u(0)=0,&x\in\mathbb{R}^{N}\setminus\{0\},\\ u\geq 0,&x\in\mathbb{R}^{N}\setminus\{0\},\end{cases}$ (1)
(where $\lambda>0$ , $2<\alpha<2p$ , $p>0$ are constants).
We mainly focus on the existence of asymptotically radially symmetric solutions and their asymptotic behavior near the rupture point with the influence of the Hénon-type term (the $|x|^{\alpha}$ term) of (1).
(Definition of asymptotic radial rupture solution: let $u(x)=u(r,\theta)$ be a nonnegative rupture solution of (1). If there exists $\beta$ such that $\lim_{r\to 0^{+}}r^{\beta}u(r,\theta)=f$ in $C^{2}(S^{N-1})$ for some constant $f$, then we say that $u(x)$ is asymptotic radial rupture solution near the origin.)
For the equation $\Delta u=u^{-p}$ with $p>0$ in $\mathbb{R}^{N}$ $(N>2)$, it has been proved in [6] that for any given $a>0$ there exists a unique radially symmetric solution $u=u(r)$ satisfying $u(0)=a$ and $u(r)\to\infty$ as $r\to\infty$.For the specfic case of $N=2$ and $F=0$ in (1).[11] investigated the isotropic and anisotropic behaviors of rupture solutions in different parameter regions, also the author studied global solutions on $R^{2}\setminus{0}$. For specfic case of $N=2$ in (1), and for different values of $\alpha$, [8] used phase-plane analysis together with the Lojasiewicz–Simon inequality to classify the asymptotic behavior of rupture solutions at the origin in the cases of asymptotically isotropic and asymptotically anisotropic profiles. Moreover, for solutions that are asymptotically isotropic at the origin, [8] obtained a more precise description and derived the first two terms in the asymptotic expansion of rupture solutions. For the specfic case $F=0$ in (1), [12]told that $u(|x|)\equiv\left(\frac{1}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}+N-2\right)}\right)^{\frac{1}{p+1}}.$ in radial cases.
For the asymptotic expansion of solutions near the fixed point. [4] use spheric harmonics to derive the asymptotic expansion near an isolated singularity for the Yamabe equation $\Delta u=\tfrac{1}{4}n(n-2)u^{\frac{n+2}{n-2}}.$ More recently, [9] also use spheric harmonics to analyze multi-term asymptotic expansions of the steady thin-film-type equation $\Delta u=u^{-p}-q,(x\in B_{R}\setminus\{0\}\quad p,q\in\mathbb{R},\ p>0)$ near the rupture point (the origin).
Although there have been some results on the existence of rupture solutions to MEMS-type elliptic equations and on their asymptotic behavior near the rupture point, to the best of our knowledge no corresponding results are available for the equation with the Hénon-type term considered here. In this paper we perform a detailed case-by-case analysis and establish the theoretical framework and results in this direction, which yields richer results. Now we give the main thereom
Theorem $1.$
For the elliptic MEMS equation (1) with $N\geq 2$, there exists at least one radial solution near the origin satisfying
$u(r)=\Lambda r^{\frac{\alpha+2}{p+1}}+\sum_{i=1}^{\infty}d_{i}r^{\frac{(2p-\alpha)i+(\alpha+2)}{p+1}}=u_{s}(r)\left(1+O\left(r^{\frac{2p-\alpha}{p+1}}\right)\right)\qquad\text{as }r\rightarrow 0^{+},$
where $\mathrm{u}_{s}(r)=\Lambda r^{\frac{\alpha+2}{p+1}}$ and $\Lambda=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}+N-2\right)}\right)^{\frac{1}{p+1}}.$
Remark. For the case of $N=1$, there exists at least one solution near the origin satisfying
$u(x)=\Lambda r^{\frac{\alpha+2}{p+1}}+\sum_{i=1}^{\infty}d_{i}x^{\frac{(2p-\alpha)i+(\alpha+2)}{p+1}}=u_{s}(x)\left(1+O\left(x^{\frac{2p-\alpha}{p+1}}\right)\right)\qquad\text{as }x\rightarrow 0^{+},$
where $\mathrm{u}_{s}(x)=\Lambda r^{\frac{\alpha+2}{p+1}}$ and $\Lambda=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}-1\right)}\right)^{\frac{1}{p+1}}.$, since the arguments for radial solution in the case $N\geq 2$ is completely similar to these for $N=1$, we shall omit the case $N=1$.
Theorem 1.2.
For the elliptic MEMS equation (1), there exist infinitely many non-radial positive asymptotic radial rupture solutions satisfying
$u(x)=\Lambda|x|^{\frac{\alpha+2}{p+1}}+C_{10}\left(\frac{x}{|x|}\right)|x|^{u_{1}}+\sum_{j=2}^{\infty}\sum_{i=0}^{i-1}C_{ji}\left(\frac{x}{|x|}\right)(\ln|x|)^{i}|x|^{u_{j}}=u_{s}(|x|)\left(1+O\left(|x|^{u_{1}}\right)\right)$
as $|x|\rightarrow 0^{+}$,
here $\mathrm{u}_{s}(|x|)=\Lambda|x|^{\frac{\alpha+2}{p+1}}$, $\Lambda$ is the same as above, $\left\{\mu_{j}\right\}_{j\geq 1}$ is a strictly increasing sequence of positive numbers diverging to $\infty$. Where
[…]
in the case $F\neq 0$.
[…]
in the case $F=0$.
And $\delta^{(k)}=\frac{-1}{2}(p+1)(N+2)+\frac{1}{2}\sqrt{(N-2)^{2}(p+1)^{2}+4(p+1)k(N-2+k)},$
$\sigma_{1}^{(k)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{1}{2}\sqrt{\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}+4k(N-2+k)-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}.$
Let $t=\ln|x|$ $\theta=\dfrac{x}{|x|}$, then
$\Delta u=u_{rr}+\frac{N-1}{r}u_{r}+\frac{1}{r^{2}}\Delta_{\theta}u=\frac{\lambda|x|^{\alpha}}{u^{p}}+F$ (2)
Let $\dfrac{\alpha+2}{p+1}=q$, $z(t,\theta)=r^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$,    where $\Lambda=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}+N-2\right)}\right)^{\frac{1}{p+1}}.$. then we can obtain:
$z_{tt}+\left(N-2+2q\right)z_{t}+(\alpha+2)(N-2+q)z+\Delta{}_{\theta}z=\frac{\lambda}{(z+\Lambda)^{p}}-\frac{\lambda}{\Lambda^{p}}+\frac{\lambda pz}{\Lambda^{p+1}}+Fr^{pq-\alpha}$
Set $f(z)=\dfrac{\lambda}{(z+\Lambda)^{p}}-\dfrac{\lambda}{\Lambda^{p}}+\dfrac{\lambda pz}{\Lambda^{p+1}}$, so
$z_{tt}+\left(N-2+2q\right)z_{t}+(\alpha+2)(N-2+q)z+\Delta{}_{\theta}z=f(z)+Fe^{\frac{2p-\alpha}{p+1}t}$ (3)
2 asymptotic behavior for radial solution
In this section, we give the arbitrary order asymptotic expansion of radial case.
Theorem 2.1.
Suppose that $u(x)$ is a radially symmetric solution of (1) with $F\neq 0$ and $\alpha<2p$. Assume it exists $\varepsilon>0$ such that
$u(r)=u_{s}(r)\left(1+O\left(r^{\varepsilon}\right)\right),~~as~r\rightarrow 0^{+},$
where $\mathrm{u}_{s}(r)=\Lambda r^{\frac{\alpha+2}{p+1}}$ and
$\Lambda=\left[\frac{\lambda(\alpha+2)}{p+1}\left(N-2+\frac{\alpha+2}{p+1}\right)\right]^{\frac{1}{p+1}}.$
Set
$z(t)=r^{-\frac{\alpha+2}{p+1}}u(r)-\Lambda,\quad t=\ln r.$
Then for any $k\gg 1$ we have
$z(t)=\sum_{\ell=1}^{k}c_{\ell}e^{\ell\rho t}+O\left(e^{(k+1)\rho t}\right),$
where $c_{\ell}$ are constants and $\rho=\frac{2p-\alpha}{p+1}$.
Proof. First, we have
$\left\{\begin{array}[]{l}\displaystyle z_{tt}+\left(N-2+2\frac{\alpha+2}{p+1}\right)z_{t}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)z=f(z)+Fe^{\frac{2p-\alpha}{p+1}t},\quad t\in(-\infty,0),\\[6.0pt] \displaystyle z(t)=O\left(e^{\varepsilon t}\right)\rightarrow 0\quad\text{as }t\rightarrow-\infty,\end{array}\right$.
where
$f(z)=\lambda(\Lambda+z)^{-p}-\lambda\Lambda^{-p}+\lambda pz\Lambda^{-(p+1)}=O(z^{2})\quad~~as~z(t)\rightarrow 0$
The associated homogeneous ordinary differential equation is
$z_{tt}+\left(N-2+2\frac{\alpha+2}{p+1}\right)z_{t}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)z=0,$
whose characteristic equation is
$\sigma^{2}+\left(N-2+2\frac{\alpha+2}{p+1}\right)\sigma+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)=0,$
it admits two roots, denoted by $\sigma_{1}^{(0)}$ and $\sigma_{2}^{(0)}$.
case1. $\alpha>\frac{(p+1)(N-2)}{2}(\sqrt{1+\frac{1}{p}}-1)-2$
$\left\{\begin{aligned} \sigma_{1}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{i}{2}\sqrt{4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-\left(N-2+2\frac{a+2}{p+1}\right)^{2}},\\ \sigma_{2}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)-\frac{i}{2}\sqrt{4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-\left(N-2+2\frac{a+2}{p+1}\right)^{2}},\end{aligned}\right$.
case2. $\alpha=\frac{(p+1)(N-2)}{2}(\sqrt{1+\frac{1}{p}}-1)-2$
$\sigma_{1}^{(0)}=\sigma_{2}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)<0;$
case3. $2<\alpha<\frac{(p+1)(N-2)}{2}(\sqrt{1+\frac{1}{p}}-1)-2$
$\left\{\begin{aligned} \sigma_{1}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{1}{2}\sqrt{\left(N-2+2\frac{a+2}{p+1}\right)^{2}-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}<0,\\ \sigma_{2}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)-\frac{1}{2}\sqrt{\left(N-2+2\frac{a+2}{p+1}\right)^{2}-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}<0,\end{aligned}\right$.
We now prove
$z(t)=O\left(e^{\frac{2p-\alpha}{p+1}t}\right)\qquad\text{as }t\to$-$\infty$
We only prove case3 and other cases are similar.By ordinary equation theory,for $T\ll-1$ and $t\in(-\infty,T)$,
$z(t)=$ $A_{1}e^{\sigma_{1}^{(0)}t}+A_{2}e^{\sigma_{2}^{(0)}t}+B_{1}e^{\sigma_{1}^{(0)}t}\int_{-\infty}^{t}e^{-\sigma_{1}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}s}\right]ds$
$+B_{2}e^{\sigma_{2}^{(0)}t}\int_{-\infty}^{t}e^{-\sigma_{2}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}s}\right]ds,$
[…]
We also know
$f(z)=O(z^{2})=O(e^{2\epsilon t}).$
on the other hand, $z(t)\to 0$ as $t\to$-$\infty$ , so $A_{1}=A_{2}=0$,
$z(t)=B_{1}e^{\sigma_{1}^{(0)}}t\int_{-\infty}^{t}e^{-\sigma_{1}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}t}\right]ds+B_{2}e^{\sigma_{2}^{(0)}}t\int_{-\infty}^{t}e^{-\sigma_{2}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}t}\right]ds$ (4)
[…]
For(ii), $f(z)=O\left(e^{\frac{2p-\alpha}{p+1}t}\right)$,it can be proved by(4).
For(i), $f(z)=O(e^{2\varepsilon t})$,we know from (4) that
$z(t)=O(e^{2\varepsilon t}).$ (5)
put(5)into $f(z(t))+Fe^{\frac{2p-\alpha}{p+1}t}$ we find
$f(z(t))+Fe^{\frac{2p-\alpha}{p+1}t}=O\left(e^{\min\left\{\frac{2p-\alpha}{p+1},4\varepsilon\right\}t}\right),$
we get by(4) that:
$z(t)=O\left(e^{\min\left\{\frac{2p-\alpha}{p+1},4\varepsilon\right\}t}\right),$ (6)
[…]
set $\rho={\frac{2p-\alpha}{p+1}}$,
$f(z)$ $=\frac{\lambda\Lambda^{p+1}\left(1-\left(1+\frac{z}{\Lambda}\right)^{p}+p\frac{z}{\Lambda}\left(1+\frac{z}{\Lambda}\right)^{p}\right)}{(1+z)^{p}\Lambda^{p+1}}$
$=d_{2}z^{2}+d_{3}z^{3}+\cdots+d_{n}z^{n}+\cdots$
$=d_{2}e^{2\rho t}+d_{3}e^{3\rho t}+\cdots+d_{n}e^{n\rho t}+\cdots$
put in into (4) we get that for any $k\gg 1$ and $t\in(-\infty,-1]$
$z(t)=\sum_{\ell=1}^{k}c_{\ell}e^{l\rho t}+O\left(e^{(k+1)\rho t}\right).$
Therefore the prove of Theorem2.1 is completed.
3 asymptotic behavior for nonradial with asymptotic radial solution
In this sction, we give the arbitrary order asymptotic expansion of the case
Theorem 3.1.
Assume $u\in C^{2}(B\backslash\{0\})$ is a positive rupture solution and exists $\epsilon>0$ satisfying $u(x)=u_{s}(|x|)\left(1+O\left(|x|^{\epsilon}\right)\right)$. Define $z(t,\theta)=r^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$,( $\Lambda$ is same as above), $\mathrm{t}=\ln|x|$, then it exists a positive sequence $\left\{\mu_{j}\right\}_{j\geq 1}$ strictly increasing to $\infty$.
When $F=0$, we have
$\mu_{1}=\begin{cases}\sigma_{1}^{(1)},&-2<\alpha<\delta^{(1)},\\[2.0pt] \sigma_{1}^{(k)},&\delta^{(k-1)}<\alpha<\delta^{(k)},\quad k=2,3,\cdots.\end{cases}$
When $F\neq 0$.
${u_{1}}=\left\{\begin{array}[]{lll}{\sigma_{1}^{(1)}}&-2<\alpha<\delta^{(1)}~and~\alpha<2p-(p+1)\sigma_{1}^{(1)}\\[2.0pt] {\sigma_{1}^{(k)}}&\delta^{(k-1)}<\alpha<\delta^{(k)}~and~\alpha<2p-(p+1)\sigma_{1}^{(k)}\\ \displaystyle\frac{2p-\alpha}{p+1}&\text{if $\delta^{(k-1)}<\alpha<\delta^{(k)}$ and $\alpha\geq 2p-(p+1)\sigma_{1}^{(k)}$,}\end{array}\right$.
For any positive interger $n\gg 1$ and $(t,\theta)\in(-\infty,-1)\times S^{N-1}$
$z(t,\theta)=\sum_{j=1}^{n}\sum_{l=0}^{j-1}c_{jl}(\theta)t^{l}e^{\mu_{j}t}+O\left(t^{n}e^{\mu_{n+1}t}\right)$
[…]
$\Delta_{S^{N-1}}Q(\theta)=\lambda_{i}Q(\theta)$
proof. It’s easy to see $z(t,\theta)=O(e^{\epsilon t})$ as $t\to$-$\infty$, and it satisfies the equation
$z_{tt}+\left(N-2+2\frac{\alpha+2}{p+1}\right)z_{t}+\Delta_{S^{N-1}}z+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)z=f(z)+Fe^{\frac{2p-\alpha}{p+1}t},\quad(t,\theta)\in(-\infty,0)\times S^{N-1},$
where $f(z)=\lambda(\Lambda+z)^{-p}-\lambda\Lambda^{-p}+\lambda pz\Lambda^{-(p+1)}=O(z^{2}).$
Considering linearization operator
$\mathcal{L}=\frac{\partial^{2}}{\partial t^{2}}+\left(N-2+2\frac{\alpha+2}{p+1}\right)\frac{\partial}{\partial t}+\Delta_{S^{N-1}}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right).$ (7)
the operator $\mathcal{L}$ can be divided into infinite partial operators
$\mathcal{L}_{k}=\frac{d^{2}}{dt^{2}}+\left(N-2++2\frac{\alpha+2}{p+1}\right)\frac{d}{dt}-\lambda_{k}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right),$ (8)
for $k=0,1,2,\ldots$, $\lambda_{k}$is the k-th eigenvalues of the eigenvalue problem
$\Delta_{S^{N-1}}Q=\lambda Q$
$\lambda_{k}=k(N-2+k)$ and
$m_{k}=\frac{(N-2+2k)(N-3+k)!}{k!(N-2)!}.$
We define $\{Q_{1}^{k}(\theta),\ldots,Q_{m_{k}}^{k}(\theta)\}$ is a set of orthonormal bases corresponding to the feature space of $\lambda_{k}$
The characteristic equation corresponding to (8) is
$\sigma^{2}+\left(N-2+2\frac{\alpha+2}{p+1}\right)\sigma+\left[(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-k(N-2+k)\right]=0.$
Its two roots are
$\sigma_{1}^{(k)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{1}{2}\sqrt{\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}+4k(N-2+k)-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}.$
$\sigma_{2}^{(k)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)-\frac{1}{2}\sqrt{\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}+4k(N-2+k)-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}.$
Fix $k=k_{0}$, so
$\sigma_{1}^{(k_{0}+1)}>\sigma_{1}^{(k_{0})}>\sigma_{1}^{(1)}>0,\quad\sigma_{2}^{(k_{0}+1)}<\sigma_{2}^{(k_{0})}<\sigma_{2}^{(1)}<0.$
where $2<\alpha<\delta^{(1)}$ ($k=k_{0}\geq 2$).
$\sigma_{1}^{(k_{0}+1)}>\sigma_{1}^{(k_{0})}>0>\sigma_{1}^{(k_{0}-1)},\quad\sigma_{2}^{(k_{0}+1)}<\sigma_{2}^{(k_{0})}<\sigma_{2}^{(1)}<0.$
where $\delta^{(k_{0}-1)}<\alpha<\delta^{(k_{0})}$ ($k=k_{0}\geq 2$)
Remark: In fact,It will be possibile that $\sigma_{1}^{(k)}$ and $\sigma_{2}^{(k)}$ are imaginary number when k is samll enough, but it will be similar with the proof below because such $\sigma^{(k)}$ is finite and $\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)<0$.
Lemma 3.2.
For $N\geq 2$, $p>0$. Assume $u$ is a positive rupture solution of (1) satisfying $u(x)=u_{s}(|x|)\left(1+O\left(|x|^{\epsilon}\right)\right)$ with
$\left\{\begin{array}[]{lll}{0<\epsilon\leq\sigma_{1}^{(1)},\qquad\text{as }-2<\alpha<\delta^{(1)}~~and~~\alpha<2p-(p+1)\sigma_{1}^{(1)}}\\ {0<\epsilon\leq\sigma_{1}^{(k)},\qquad\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}~~and~~\alpha<2p-(p+1)\sigma_{1}^{(k)}}\\ {0<\epsilon\leq\frac{2p-\alpha}{p+1},\qquad\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}~~and~~\alpha\geq 2p-(p+1)\sigma_{1}^{(k)}}\end{array}\right$.
If we define $z(t,\theta)=|x|^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$ ($t=\ln r$), then $z(t,\theta)=\mathcal{O}(e^{\varepsilon t})$ for t $\in(-\infty,1)$ , and for the case $F\neq 0$,
[…]
For the case $F=0$,
$\max_{S^{N-1}}|z(t,\theta)|\leq\left\{\begin{array}[]{ll}{\sigma_{1}^{(1)}}&\text{if }-2<\alpha<\delta^{(1)},\\ {\sigma_{1}^{(k)}}&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{array}\right$.
Let
$\overline{z}(t)=\frac{\int_{\mathbb{S}^{N-1}}z(t,\theta)\,d\theta}{|\mathbb{S}^{N-1}|},\quad w(t,\theta)=z(t,\theta)-\overline{z}(t).$
Since the Laplace operator on the sphere has zero integral over $\mathbb{S}^{N-1}$, we observe that $w$ satisfies the following equation:
$w_{tt}+\left(N-2+2\frac{\alpha+2}{p+1}\right)w_{t}+\Delta_{\mathbb{S}^{N-1}}w+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)w=f(z)-\overline{f(z)},$ (9)
It is clear that $w(t,\theta)=\mathcal{O}(e^{\varepsilon t})~~as~t\to$-$\infty$.
A direct computation shows that
$f(z)-\overline{f(z)}=f^{\prime}(\xi)\,w-f^{\prime}(\xi)\,\overline{w},$ (10)
where $\xi$ lies between $z$ and $\overline{z}$, and satisfies $\xi=\mathcal{O}(e^{\varepsilon t})~~as~t\to$-$\infty$.
Furthermore, we obtain
$f^{\prime}(\xi)=-\lambda p\left(\Lambda+\xi\right)^{-(p+1)}+\lambda p\Lambda^{-(p+1)}=\mathcal{O}(e^{\varepsilon t}).~~~~~as~~t\to$-$\infty$
We first derive an estimate for $w$.
Lemma 3.3.
Assume that $w(t,\theta)$ is a solution of (9). Then
$\max_{S^{N-1}}|w(t,\theta)|\leq\left\{\begin{array}[]{ll}\sigma_{1}^{(1)},&\text{if }-2<\alpha<\delta^{(1)},\\[2.84526pt] \sigma_{1}^{(k)},&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{array}\right$.
we focus on the case $\delta^{(k-1)}<\alpha<\delta^{(k)}$, the remaining case is similiar and eaiser.
For notational convenience, we fix $k_{0}$ such that $\delta^{(k_{0}-1)}<\alpha<\delta^{(k_{0})}$.
We expand $w$ as
$w(t,\theta)=\sum_{k=1}^{\infty}\sum_{j=0}^{m_{k}}w_{j}^{k}(t)\,Q_{j}^{k}(\theta).$
Since the first item of $z(t,\theta)-\overline{z}(t)$ vanishes, we have $m_{0}=1$ and $w_{1}^{(0)}(t)\equiv 0$. $w_{j}^{(k)}(t)$ satisfies the ODE
$\left(w_{j}^{(k)}\right)^{\prime\prime}+\left(N-2+2\frac{\alpha+2}{p+1}\right)\left(w_{j}^{(k)}\right)^{\prime}+\left[(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-\lambda_{k}\right]w_{j}^{(k)}=g_{j}^{(k)}(t),$ (11)
where
$g_{j}^{(k)}(t)$ $=\int_{\mathbb{S}^{N-1}}\left(f(z(t,\theta))-\overline{f(z(t,\theta))}\right)Q_{j}^{(k)}(\theta)\,d\theta$
$=\int_{\mathbb{S}^{N-1}}\left(f(z(t,\theta))-f(\overline{z}(t))\right)Q_{j}^{(k)}(\theta)\,d\theta-\int_{\mathbb{S}^{N-1}}\overline{\left(f(z(t,\theta))-f(\overline{z}(t))\right)}Q_{j}^{(k)}(\theta)\,d\theta$
$=\int_{\mathbb{S}^{N-1}}\left(f(z(t,\theta))-f(\overline{z}(t))\right)Q_{j}^{(k)}(\theta)\,d\theta.$
For $k\geq 1$, we observe that
$\|w\|_{L^{2}(\mathbb{S}^{N-1})}^{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{(k)}(t)\right)^{2},\quad|f(z)-f(\overline{z})|=\sum_{k=1}^{\infty}\sum_{j=0}^{m_{k}}\left(g_{j}^{(k)}(t)\right)^{2}$
Since $f(z)-f(\overline{z})=f^{\prime}(\xi)\,w$ and $f^{\prime}(\xi)=\mathcal{O}(e^{\varepsilon t})$, it follows that
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(g_{j}^{(k)}(t)\right)^{2}=\mathcal{O}(e^{2\varepsilon t})\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{k}^{(j)}(t)\right)^{2}.$ (12)
On the other hand, applying (11), we obtain for $T\ll-1$ and $t<T$,
$w_{j}^{k}(t)=A_{j,1}^{k}e^{\sigma_{1}^{(k)}t}+A_{j,2}^{k}e^{\sigma_{2}^{(k)}t}+B_{j,1}^{k}\int_{t}^{T}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)\,ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)\,ds,$ (13)
where $\left|B_{j,1}^{k}\right|$ and $\left|B_{j,2}^{k}\right|$ are positive constants.
For $k\geq k_{0}$, we have $\sigma_{1}^{(k)}>0$ and $\sigma_{2}^{(k)}<0$. Since $w_{j}^{k}(t)\to 0$ as $t\to$-$\infty$, it follows that $A_{j,2}^{k}=0$. Consequently,
$w_{j}^{k}(T)=A_{j,1}^{k}e^{\sigma_{1}^{(k)}T}-B_{j,2}^{k}\int_{-\infty}^{T}e^{\sigma_{2}^{(k)}(T-s)}g_{j}^{k}(s)\,ds,$
where $A_{j,1}^{k}=O\left(e^{-\sigma_{1}^{(k)}T}\right)$.
Therefore,
$w_{j}^{k}(t)=O\left(e^{\sigma_{1}^{(k)}(t-T)}\right)+B_{j,1}^{k}\int_{t}^{T}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)\,ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)\,ds.$ (12)
We may choose $\delta>0$ sufficiently small such that
$\left[w_{j}^{k}(t)\right]^{2}\leq$ $O\left(e^{2\sigma_{1}^{(k)}(t-T)}\right)+4(B_{j,1}^{k})^{2}\left(\int_{t}^{T}e^{\delta(t-s)}\,ds\right)\left(\int_{t}^{T}e^{(2\sigma_{1}^{(k)}-\delta)(t-s)}(g_{j}^{k}(s))^{2}\,ds\right)$
$\quad+4(B_{j,2}^{k})^{2}\left(\int_{-\infty}^{t}e^{-\delta(t-s)}\,ds\right)\left(\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k)}+\delta)(t-s)}(g_{j}^{k}(s))^{2}\,ds\right)$
$\leq$ $Ce^{2\sigma_{1}^{(k)}(t-T)}+C_{\delta}\int_{t}^{T}e^{(2\sigma_{1}^{(2)}-\delta)(t-s)}(g_{j}^{k}(s))^{2}\,ds+C_{\delta}\int_{-\infty}^{t}e^{(2\sigma_{2}^{(2)}+\delta)(t-s)}(g_{j}^{k}(s))^{2}\,ds,$
where $C>0$ and $C_{\delta}>0$.
For $k<k_{0}$,
$w_{j}^{k}(t)=$ $A_{j,1}^{k}e^{\sigma_{1}^{(k)}t}+A_{j,2}^{k}e^{\sigma_{2}^{(k)}t}-B_{j,1}^{k}\int_{-\infty}^{t}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)ds$
In this case,
$\left|B_{j,1}^{k}\right|=\left|B_{j,2}^{k}\right|=\left|\frac{1}{\sigma_{2}^{(k)}-\sigma_{1}^{(k)}}\right|.$
Since $\sigma_{1}^{(k)}<0$ , $\sigma_{2}^{(k)}<0$ and $w_{j}^{k}(t)\to 0$ as $t\to$-$\infty$, we get that $A_{j,1}^{k}=A_{j,2}^{k}=0$. Hence,
$w_{j}^{k}(t)=-B_{j,1}^{k}\int_{-\infty}^{t}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)\,ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)\,ds.$ (14)
Moreover,
$\left(w_{j}^{k}(t)\right)^{2}=O\left(e^{4\varepsilon t}\right).$ (15)
Observe that
$\left(g_{j}^{k}(t)\right)^{2}\leq C\,\|f^{\prime}(\xi)w\|_{L^{2}(S^{N-1})}^{2}\leq Ce^{4\varepsilon t}.$
Therefore,
$\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}\leq$ $\;C\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}e^{2\sigma_{1}^{(k)}(t-T)}+C_{\delta}\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(g_{j}^{k}(s)\right)^{2}\,ds$
$\;+C_{\delta}\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(g_{j}^{k}(s)\right)^{2}\,ds$
$\leq$ $\;C\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}e^{2\sigma_{1}^{(k)}(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{4\varepsilon s}\,ds$
$\;+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(s)\right)^{2}\,ds$
$\;+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{4\varepsilon s}\,ds$
$\;+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(s)\right)^{2}\,ds.$
Note that
$\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(g_{j}^{k}(t)\right)^{2}\leq\|f^{\prime}(\xi)w\|_{L^{2}(S^{N-1})}^{2}-\sum_{k=1}^{k_{0}-1}\sum_{j=1}^{m_{k}}\left(g_{j}^{k}(t)\right)^{2},$
and
$\|f^{\prime}(\xi)w\|_{L^{2}(S^{N-1})}^{2}=O\left(e^{2\varepsilon t}\right)\|w\|_{L^{2}(S^{N-1})}^{2}=O\left(e^{2\varepsilon t}\right)\left[\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}+\sum_{k=1}^{k_{0}-1}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}\right]$ (3.1)
Since
$\lim_{k\to\infty}\frac{m_{k+1}e^{2(\sigma_{1}^{(k+1)}-\sigma_{1}^{(k_{0})})(t-T)}}{m_{k}e^{2(\sigma_{1}^{(k)}-\sigma_{1}^{(k_{0})})(t-T)}}=\lim_{k\to\infty}\left[\frac{m_{k+1}}{m_{k}}e^{2(\sigma_{1}^{(k+1)}-\sigma_{1}^{(k)})(t-T)}\right]=e^{(t-T)}<1,$
we obtain
$\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}e^{2\sigma_{1}^{(k)}(t-T)}=\sum_{k=k_{0}}^{\infty}m_{k}e^{2\sigma_{1}^{(k)}(t-T)}=O\!\left(e^{2\sigma_{1}^{(k_{0})}(t-T)}\right),$
Let
$W(t)]^{2}=\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}.$
Then
${}^{2}\leq$ $\;Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+Ce^{4\varepsilon t}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{4\varepsilon s}\,ds$
$\quad+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.$
Next we consider two cases: (i) $4\varepsilon\geq 2\sigma_{1}^{(2)}-\delta$, (ii) $4\varepsilon<2\sigma_{1}^{(2)}-\delta$.
For the first case, we first assume $4\varepsilon>2\sigma_{1}^{(2)}-\delta$. We have
${}^{2}\leq$ $Ce^{(2\sigma_{1}^{(2)}-\delta)(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(2)}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds$ (16)
$\quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.$
Define
$K_{1}(t)=\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}[W(s)]^{2}\,ds,\qquad K_{2}(t)=\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}[W(s)]^{2}\,ds.$
For $T$ sufficiently small, we compute that
$\left(K_{2}-K_{1}\right)^{\prime}(t)$ $=(2\sigma_{2}^{(k_{0})}+\delta)K_{2}(t)-(2\sigma_{1}^{(k_{0})}-\delta)K_{1}(t)+2[W(t)]^{2}$
$\leq(2\sigma_{2}^{(k_{0})}+\delta)K_{2}(t)-(2\sigma_{1}^{(k_{0})}-\delta)K_{1}(t)+Ce^{2\varepsilon T}\left(K_{1}(t)+K_{2}(t)\right)+Ce^{(2\sigma_{1}^{(k_{0})}-\delta)(t-T)}$
$\leq Ce^{(2\sigma_{1}^{(k_{0})}-\delta)(t-T)}.$
Note $\sigma_{2}^{(k_{0})}<0$ and $\sigma_{1}^{(k_{0})}>0$, and choose $\delta>0$ sufficiently small. Since $K_{1}(t)\to 0$ and $K_{2}(t)\to 0$ as $t\to$-$\infty$, we obtain for all $t<T$,
$K_{2}(t)\leq K_{1}(t)+Ce^{(2\sigma_{1}^{(k_{0})}-\delta)t}.$ (17)
Substituting (17) into (16), we get at
$W(t)]^{2}\leq Ce^{(2\sigma_{1}^{(k_{0})}-\delta)t}+Ce^{2\varepsilon T}K_{1}(t).$ (18)
From (18), it follows that
$K_{1}(t)\leq\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}\left[Ce^{(2\sigma_{1}^{(k_{0})}-\delta)s}+Ce^{2\varepsilon T}K_{1}(s)\right]ds.$
Therefore,
$e^{-2(\sigma_{1}^{(k_{0})}-\delta)t}\,K_{1}(t)\leq C(T-t)+Ce^{2\varepsilon T}\int_{t}^{T}K_{1}(s)\,e^{-(2\sigma_{1}^{(k_{0})}-\delta)s}\,ds.$
Let $F_{1}(t)=\int_{t}^{T}e^{-(2\sigma_{1}^{(k_{0})}-\delta)s}\,ds$. Then
$F_{1}^{\prime}(t)\leq C(T-t)+Ce^{2\varepsilon T}F_{1}(t).$
Setting $\mu=Ce^{2\varepsilon T}$, we obtain
$(e^{\mu t}F_{1}(t))^{\prime}\leq C(T-t)e^{\mu t}.$
Integrating over $(t,T)$ yields
$F_{1}(t)\leq\frac{C}{\mu^{2}}e^{-\mu(t-T)}.$
Hence,
$K_{1}(t)$ $\leq Ce^{(2\sigma_{1}^{(k_{0})}-\delta)t}(T-t)+Ce^{2\mu}e^{(2\sigma_{1}^{(k_{0})}-\delta)t}F_{1}(t)$
$\leq Ce^{(2\sigma_{1}^{(k_{0})}-\delta)t}(T-t)+\frac{C}{\mu}e^{(2\sigma_{1}^{(k_{0})}-\delta-\mu)t}e^{\mu T}.$
Therefore,
$K_{1}(t)=O\left(e^{(2\sigma_{1}^{(k_{0})}-\delta-\mu)t}\right).$
Consequently,
$W(t)]^{2}\leq C\mu\,e^{(2\sigma_{1}^{(k_{0})}-\delta-\mu)t}.$
Meanwhile, by (15) and the assumption $4\varepsilon\geq 2\sigma_{1}^{(k_{0})}-\delta$, we also obtain
$\left(w_{j}^{k}(t)\right)^{2}=O\left(e^{(2\sigma_{1}^{(2)}-\delta)t}\right),\qquad k=1,2,\ldots,k_{0}-1.$
Hence, for all $t<T$,
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}=O\left(e^{(2\sigma_{1}^{(k_{0})}-\delta-\mu)t}\right),$
therefore
$\|w\|_{L^{2}(S^{N-1})}=O\left(e^{\left(\sigma_{1}^{(k_{0})}-\frac{\delta}{2}-\frac{\mu}{2}\right)t}\right).$ (19)
$\max_{S^{N-1}}|w(t,\theta)|\leq Me^{\left(\sigma_{1}^{(k_{0})}-\frac{\delta}{2}-\frac{\mu}{2}\right)t},\qquad\text{for all }t\in(-\infty,-1].$ (20)
We establish (20) only for $t\in(-\infty,T_{*})$, ($T_{*}\leq T$). The remaining part can be proved directly from the continuity of $w$. Define
$h(r,\theta)=w(t,\theta),\qquad r=e^{t}.$
Then $h(r,\theta)$ satisfies the equation
$\Delta h+\frac{b_{1}\,x\cdot\nabla h}{r^{2}}+\frac{b_{2}h}{r^{2}}-\frac{f^{\prime}(\xi)h-\overline{f^{\prime}(\xi)h}}{r^{2}}=0,\qquad\text{in }B_{R_{*}}\setminus\{0\},$ (21)
where
$b_{1}=2\frac{\alpha+2}{p+1},\qquad b_{2}=2\left(N-2+\frac{\alpha+2}{p+1}\right),\qquad R_{*}=e^{T_{*}}.$
For any $x_{0}\in B_{R_{*}}\setminus\{0\}$, denote $r_{0}=|x_{0}|$ and set $\Omega=B_{r_{0}/2}(x_{0})$. We regard equation (21) as the linear equation appearing in Lemma 5.1 of reference [3], where
$k=k_{1}=1,\qquad h(x)\equiv 0,\qquad|b|=\frac{4b_{1}^{2}}{r_{0}^{2}},\qquad|c|=\frac{Q}{r_{0}^{2}},$
Here $Q=Q(h)>0$. Thus, by applying Lemma 5.1 in reference [3] with $k_{2}=Q/r_{0}^{2}$, and combining (19) with an argument similar to the proof of Theorem 5.1 in reference [3], we conclude that there exists a positive constant
$M=M\!\left(\frac{k_{1}}{k},\,k_{2}r_{0}^{2}\right)=M(Q)=M(h),$
independent of $r_{0}$, such that
$\sup_{x\in B_{r_{0}/4}(x_{0})}|h(x)|\leq M\,r_{0}^{\sigma_{1}^{(k_{0})}-\frac{\delta}{2}-\frac{\mu}{2}}.$
In particular,
$|h(x_{0})|\leq M\,r_{0}^{\sigma_{1}^{(k_{0})}-\frac{\delta}{2}-\frac{\mu}{2}},\qquad\max_{|x|=r}|h(x)|\leq M\,r^{\sigma_{1}^{(k_{0})}-\frac{\delta}{2}-\frac{\mu}{2}}.$
We obtain from estimate (20) that
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(g_{j}^{k}(t)\right)^{2}=O\!\left(e^{(2\sigma_{1}^{(k_{0})}+2\varepsilon-\delta-\mu)t}\right).$ (22)
By (14) and (22), we know
$w_{j}^{k}(t)=O\!\left(e^{(\sigma_{1}^{(k_{0})}+\varepsilon-\frac{\delta}{2}-\frac{\mu}{2})t}\right)=O\!\left(e^{\sigma_{1}^{(k_{0})}t}\right),\qquad k\leq k_{0},$ (23)
since we may choose $\delta$ sufficiently small and $T$ sufficiently large so that $\delta<2\varepsilon-\mu$.
Moreover, by using (22) we have
$\sum_{j=1}^{m_{k_{0}}}|w_{j}^{k_{0}}(t)|\leq$ $Ce^{\sigma_{1}^{(k_{0})}t}+C\int_{t}^{T}e^{\sigma_{1}^{(k_{0})}(t-s)}\sum_{j=1}^{m_{k_{0}}}|g_{j}^{k_{0}}(s)|\,ds$
$\quad+C\int_{-\infty}^{t}e^{\sigma_{2}^{(k_{0})}(t-s)}\sum_{j=1}^{m_{k_{0}}}|g_{j}^{k_{0}}(s)|\,ds$
$\leq$ $Ce^{\sigma_{1}^{(k_{0})}t}+C\int_{t}^{T}e^{\sigma_{1}^{(k_{0})}(t-s)}e^{(\sigma_{1}^{(k_{0})}+\varepsilon-\frac{\delta}{2}-\frac{\mu}{2})s}\,ds$
$\quad+C\int_{-\infty}^{t}e^{\sigma_{2}^{(k_{0})}(t-s)}e^{(\sigma_{1}^{(k_{0})}+\varepsilon-\frac{\delta}{2}-\frac{\mu}{2})s}\,ds$
$\leq$ $Ce^{\sigma_{1}^{(k_{0})}t}.$
Since $\delta<2\varepsilon-\mu$, it follows for $t<T$ that
$\sum_{j=1}^{m_{k_{0}}}|w_{j}^{k_{0}}(t)|^{2}\leq\left[\sum_{j=1}^{m_{k_{0}}}|w_{j}^{k_{0}}(t)|\right]^{2}\leq Ce^{2\sigma_{1}^{(k_{0})}t}.$ (24)
Similarly, for $k>k_{0}$ we obtain
$\sum_{j=1}^{m_{k}}\left|w_{j}^{k}(t)\right|\leq$ $Ce^{\sigma_{1}^{(k)}t}+C\int_{t}^{T}e^{\sigma_{1}^{(k)}(t-s)}\sum_{j=1}^{m_{k}}\left|g_{j}^{2}(s)\right|ds+C\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}\sum_{j=1}^{m_{k}}\left|g_{j}^{k}(s)\right|ds$
$\leq$ $Ce^{\sigma_{1}^{(k)}t}+C\int_{t}^{T}e^{\sigma_{1}^{(k)}(t-s)}e^{\left(\sigma_{1}^{(k)}+\varepsilon-\frac{\delta}{2}-\frac{\mu}{2}\right)s}ds+C\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}e^{\left(\sigma_{1}^{(k)}+\varepsilon-\frac{\delta}{2}-\frac{\mu}{2}\right)s}ds$
$\leq$ $Ce^{\sigma_{1}^{(k)}t}.$
Hence
$\sum_{k=k_{0}+1}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}=O\!\left(e^{\sigma_{1}^{(k_{0})}t}\right).$ (25)
Since $\delta$ can be chosen sufficiently small, combining (23)–(25) yields
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}=O\!\left(e^{\sigma_{1}^{(k_{0})}t}\right).$ (26)
The case $4\varepsilon=2\sigma_{1}^{(2)}-\delta$ can be down similarly. Indeed, by enlarging $\delta$ slightly to a number $\delta^{\prime}>\delta$ such that $4\varepsilon>2\sigma_{1}^{(2)}-\delta^{\prime}$, and repeating the proof above with $\delta^{\prime}$ in stead of $\delta$, we conclude (26).
For case (ii), we have
${}^{2}\leq$ $Ce^{4\varepsilon t}+C\int_{t}^{T}e^{(2\sigma_{1}^{(2)}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds$
$\quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(2)}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.$
Since $W(t)]^{2}=O(e^{2\varepsilon t})$, it follows that
$W(t)]^{2}\leq Ce^{4\varepsilon t},\qquad t<T.$
Together with (15), this also implies
$\|w\|_{L^{2}(S^{N-1})}=O\!\left(e^{2\varepsilon t}\right).$
Thus, by an argument similar to the proof of (20), there exists a constant $M=M(w)>0$ such that
$\max_{S^{N-1}}|w(t,\theta)|\leq Me^{2\varepsilon t},\qquad t\in(-\infty,-1].$ (27)
From (14) and (27), we obtain, for $k<k_{0}$,
$w_{j}^{k}(t)=O\!\left(e^{3\varepsilon t}\right).$
Therefore, we get the inequality
$\begin{gathered}{}^{2}\leq Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{6\varepsilon s}\,ds+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds\\ \quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{6\varepsilon s}\,ds+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.\end{gathered}$ (28)
Note that:
$\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(g_{j}^{k}(t)\right)^{2}\leq Ce^{2\epsilon t}\left([W(t)]^{2}+e^{6\epsilon t}\right)$
We still consider two cases: (a) $6\varepsilon\geq 2\sigma_{1}^{(k_{0})}-\delta$, (b) $6\varepsilon<2\sigma_{1}^{(k_{0})}-\delta$.
For case (a), using inequality (28) and an argument similar in case (i), we conclude that (26) holds.
For case (b), inequality (28) implies that
$W(t)]^{2}\leq Ce^{6\varepsilon t},\qquad\text{for }t<T.$
Hence,
$\begin{gathered}{}^{2}\leq Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{8\varepsilon s}\,ds+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds\\ \quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{8\varepsilon s}\,ds+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.\end{gathered}$
We still consider into two cases: $8\varepsilon\geq 2\sigma_{1}^{(k_{0})}-\delta$ and $8\varepsilon<2\sigma_{1}^{(k_{0})}-\delta$. Repeating the same steps as before, we eventually deduce that (26) holds.
Clearly, (26) implies
$\|w\|_{L^{2}(S^{N-1})}=O\!\left(e^{\sigma_{1}^{(k_{0})}t}\right).$
Using the argument similar as in Theorem 5.1 of reference [3], we also obtain
$\max_{S^{N-1}}|w(t,\theta)|\leq Ce^{\sigma_{1}^{(k_{0})}t},\qquad t\in(-\infty,-1].$
This completes the proof of the lemma.
Lemma 3.4.
Let $\bar{z}(t)$ be defined as above. Then:
$|\bar{z}(t)|\leq\begin{cases}Ce^{\min\left\{2\sigma_{1}^{(1)},\,\frac{2p-\alpha}{p+1}\right\}t},&\text{if }-2<\alpha<\delta^{(1)},\\[2.84526pt] Ce^{\min\left\{2\sigma_{1}^{(k)},\,\frac{2p-\alpha}{p+1}\right\}t},&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{cases}$
For $F\neq 0$,
$|\bar{z}(t)|\leq\begin{cases}Ce^{2\sigma_{1}^{(1)}t},&\text{if }-2<\alpha<\delta^{(1)},\\[2.84526pt] Ce^{2\sigma_{1}^{(k)}t},&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{cases}$
For $F=0$. $C>0$ is a constant independent of $t$.
Proof. We only prove the case $F\neq 0$ with $\delta^{(k-1)}<\alpha<\delta^{(k)}$, the other cases are similar and eaiser. we only need to prove the estimate for $t$ sufficiently close to -$\infty$ and the other part can be gotten easily by the continuity of $\bar{z}(t)$,
We assume $\frac{2p-\alpha}{p+1}\leq 2\sigma_{1}^{(k)},$ while the opposite case $\frac{2p-\alpha}{p+1}>2\sigma_{1}^{(k)}$ can be treated similarly.
We know that $\bar{z}(t)$ satisfies the ODE
$\left\{\begin{array}[]{l}\bar{z}_{tt}+\left(N-2+2\frac{\alpha+2}{p+1}\right)\bar{z}_{t}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)\bar{z}=\overline{f(z)}+Fe^{\frac{2p-\alpha}{p+1}t},\quad t\in(-\infty,0),\\[2.84526pt] \bar{z}(t)=O(e^{\varepsilon t})\to 0\quad\text{as}~t\to-\infty,\end{array}\right$.
where $\varepsilon>0$ is as in Lemma 2.3.
Observe that
$\overline{f(z)}+Fe^{\frac{2p-\alpha}{p+1}t}=O\big(\overline{(\bar{z}+w)^{2}}\big)+Fe^{\frac{2p-\alpha}{p+1}t}=O\big(\overline{\bar{z}^{2}+2\bar{z}w+w^{2}}\big)+Fe^{\frac{2p-\alpha}{p+1}t}.$
By Lemma 2.4
$\overline{f(z)}+Fe^{\frac{2p-\alpha}{p+1}t}$ $=O\left(\bar{z}^{2}+2\bar{z}\,O(e^{\sigma_{1}^{(k)}t})+O(e^{2\sigma_{1}^{(k)}t})\right)+Fe^{\frac{2p-\alpha}{p+1}t}$
$=O(e^{2\varepsilon t})+O(e^{(\sigma_{1}^{(k)}+\varepsilon)t})+O\!\left(e^{\min\{2\sigma_{1}^{(k)},\,\frac{2p-\alpha}{p+1}\}t}\right)$
$=O(e^{2\varepsilon t})+O\left(e^{\frac{2p-\alpha}{p+1}t}\right).$
as $t\to$-$\infty$.
Since
$\frac{2p-\alpha}{p+1}\leq 2\sigma_{1}^{(k)},\qquad 0<\varepsilon\leq\sigma_{1}^{(k)},$
we consider two cases: (i) $2\varepsilon\geq\frac{2p-\alpha}{p+1}$, (ii) $2\varepsilon<\frac{2p-\alpha}{p+1}$.
Case (i): $2\varepsilon\geq\frac{2p-\alpha}{p+1}$
In this case,
$\overline{f(z)}+Fe^{\frac{2p-\alpha}{p+1}t}=O\left(e^{\frac{2p-\alpha}{p+1}t}\right).$
Using the representation formula for solutions of the ODE for $\bar{z}$, and following the similar argument as in the proof of Theorem 2.1, we obtain
$\bar{z}(t)=O\left(e^{\frac{2p-\alpha}{p+1}t}\right).$
This completes the proof of Lemma 2.5 in this case.
Case (ii): $2\varepsilon<\frac{2p-\alpha}{p+1}$
We have
$\overline{f(z)}+Fe^{\frac{2p-\alpha}{p+1}t}=O(e^{2\varepsilon t}).$
Using again the ODE theory for $\bar{z}$ and an argument similar to the proof of Theorem 2.1, we conclude that
$\bar{z}(t)=O(e^{2\varepsilon t}).$
By Lemma 2.4, we have
$z(t,\theta)=w(t,\theta)+\bar{z}(t)=O\!\left(e^{\sigma_{1}^{(k)}t}\right)+O\!\left(e^{2\varepsilon t}\right).$
we obtain
$z(t,\theta)=O\!\left(e^{\sigma_{1}^{(k)}t}\right),\qquad f(z)=O(z^{2})=O\!\left(e^{2\sigma_{1}^{(k)}t}\right).$
When $2\sigma_{1}^{(k)}\leq\varepsilon$.
Thus
$\overline{f(z)}+Fe^{\frac{2p-\alpha}{p+1}t}=O\!\left(e^{\frac{2p-\alpha}{p+1}t}\right).$
Using the ODE for $\bar{z}$, we obtain
$\bar{z}(t)=O\!\left(e^{\frac{2p-\alpha}{p+1}t}\right).$
Also we obtain
$z(t,\theta)=O\!\left(e^{2\varepsilon t}\right),$
When $2\varepsilon<\sigma_{1}^{(k)}$, and therefore
$\overline{f(z)}+Fe^{\frac{2p-\alpha}{p+1}t}=O\!\left(e^{4\varepsilon t}\right)+O\!\left(e^{\frac{2p-\alpha}{p+1}t}\right).$
We still consider two cases: (a) $4\varepsilon\geq\frac{2p-\alpha}{p+1}$, (b) $4\varepsilon<\frac{2p-\alpha}{p+1}$. Repeating the similar steps as above then we yield the desired estimate. We now prove Lemma 2.3. From the arguments given above, observe that when $\alpha<(\geq)\,2p-(p+1)\sigma_{1}^{(k)}$, we have $\sigma_{1}^{(k)}<(\geq)\,\frac{2p-\alpha}{p+1}$. Since $z(t,\theta)=w(t,\theta)+\bar{z}(t)$, the conclusion follows immediately.
Finally, we proceed to prove Theorem 3.1.
(i) For the case $F=0$ and $2<\alpha<\delta^{(1)}$, we set
$\rho_{1}=\sigma_{1}^{(1)},\;\rho_{2}=\sigma_{1}^{(2)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m)},\;\ldots$ (29)
For the case $F=0$ and $\delta^{(k-1)}<\alpha<\delta^{(k)}$, we set
$\rho_{1}=\sigma_{1}^{(k)},\;\rho_{2}=\sigma_{1}^{(k+1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(k+m)},\;\ldots$ (30)
(ii) For the case $F\neq 0$ and $2<\alpha<\delta^{(1)}$, if there exists some $j$ such that
$\rho_{1}^{(j)}<\frac{2p-\alpha}{p+1}<\rho_{1}^{(j+1)},$
we set
$\rho_{1}=\sigma_{1}^{(1)},\;\ldots,\;\rho_{j+1}=\frac{2p-\alpha}{p+1},\;\rho_{j+2}=\sigma_{1}^{(j+1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m)},\;\ldots$ (31)
For the case $F\neq 0$ and $2<\alpha<\delta^{(1)}$, if there exists some $j$ such that
$\sigma_{1}^{(j)}<\frac{2p-\alpha}{p+1}=\sigma_{1}^{(j+1)}\quad\text{or}\quad\sigma_{1}^{(j)}=\frac{2p-\alpha}{p+1}<\sigma_{1}^{(j+1)},$
we set
$\rho_{1}=\sigma_{1}^{(1)},\;\ldots,\;\rho_{j}=\sigma_{1}^{(j)},\;\rho_{j+1}=\sigma_{1}^{(j+1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m)},\;\ldots$ (32)
For the case $F\neq 0$ and $2<\alpha<\delta^{(1)}$, if
$\frac{2p-\alpha}{p+1}<\sigma_{1}^{(1)},$
we set
$\rho_{1}=\frac{2p-\alpha}{p+1},\;\rho_{2}=\sigma_{1}^{(1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m-1)},\;\ldots$ (33)
For the case $F\neq 0$ and $\delta^{(k-1)}<\alpha<\delta^{(k)}$, if there exists some $j\geq k$ such that
$\sigma_{1}^{(j)}<\frac{2p-\alpha}{p+1}<\sigma_{1}^{(j+1)},$
we set
$\rho_{1}=\sigma_{1}^{(k)},\;\rho_{2}=\sigma_{1}^{(k+1)},\;\ldots,\;\rho_{j-k+1}=\sigma_{1}^{(j)},\;\rho_{j-k+2}=\frac{2p-\alpha}{p+1},\;\rho_{j-k+3}=\sigma_{1}^{(j+1)},\;\ldots$ (34)
For the case $F\neq 0$ and $\delta^{(k-1)}<\alpha<\delta^{(k)}$, if there exists some $j\geq k$ such that
$\sigma_{1}^{(j)}<\frac{2p-\alpha}{p+1}=\sigma_{1}^{(j+1)}\quad\text{or}\quad\sigma_{1}^{(j)}=\frac{2p-\alpha}{p+1}<\sigma_{1}^{(j+1)},$
we set
$\rho_{1}=\sigma_{1}^{(k)},\;\rho_{2}=\sigma_{1}^{(k+1)},\;\ldots,\;\rho_{j-k+1}=\sigma_{1}^{(j)},\;\rho_{j-k+2}=\sigma_{1}^{(j+1)},\;\ldots$ (35)
For the case $F\neq 0$ and $\delta^{(k-1)}<\alpha<\delta^{(k)}$, if
$\frac{2p-\alpha}{p+1}<\sigma_{1}^{(k)},$
we set
$\rho_{1}=\frac{2p-\alpha}{p+1},\quad\rho_{2}=\sigma_{1}^{(k)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m-2+k)},\;\ldots$ (36)
The sequence $\{\rho_{i}\}_{i\geq 1}$ is strictly increasing and diverges to $\infty$ under the above assumptions,
We first prove case (29), and then indicate the proof of the remaining cases.
We start from the identity
$\mathcal{L}z=f(z),$
where
$\mathcal{L}z=z_{tt}+\left(N-2+2\frac{\alpha+2}{p+1}\right)z_{t}+\Delta_{S^{N-1}}z+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)z,$
and
$f(z)=\lambda(\Lambda+z)^{-p}-\lambda\Lambda^{-p}+\lambda pz\Lambda^{-(p+1)}.$
For $\Delta_{\theta}Q_{i}=\lambda_{i}Q_{i}$ $(i\geq 0)$, and for clarity of presentation, we write the eigenvalues with multiplicities:
$\lambda_{0}=0,\qquad\lambda_{1}=\cdots=\lambda_{n}=1,\qquad\lambda_{n+1}=2n,\quad\ldots$
We fix $\{Q_{i}\}$ as an orthonormal basis of $L^{2}(\mathbb{S}^{n-1})$.
For each fixed $i\geq 0$ and any twice differentiable function $\psi=\psi(t)$, we define
$\mathcal{L}(\psi Q_{i})=(L_{i}\psi)\,Q_{i}.$
Since $\Delta_{\theta}Q_{i}=\lambda_{i}Q_{i}$, we obtain
$L_{i}\psi=\psi_{tt}+\left(N-2+2\frac{\alpha+2}{p+1}\right)\psi_{t}-\lambda_{i}\psi+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)\psi.$
Lemma 3.5.
There exist two sequences $\{\rho_{i}\}_{i\geq 1}$ and $\{\tau_{i}\}_{i\geq 1}$, tending respectively to $\infty$ and -$\infty$, such that for every $i\geq 1$, $\operatorname{Ker}(L_{i})$ has a basis $\{e^{\rho_{i}t},\,e^{\tau_{i}t}\}$.
We now introduce the notion of an index set. Let $\{\rho_{i}\}_{i\geq 1}$ denote the sequence appearing in Lemma 3.5 (i.e., the sequence $\{\rho_{i}\}$ in (29) with multiplicities considered), which is strictly increasing and diverges to $\infty$.
We define the index set
[…]
In other words, $\mathcal{I}$ consists of all finite positive-integer linear combinations of the $\rho_{i}$. It’s possible that a given $\rho_{i}$ may itself be representable as a positive-integer linear combination of $\rho_{1},\ldots,\rho_{i-1}$.
We now give another lemma.
Lemma 3.6.
If $Q_{k}$ and $Q_{l}$ are spherical harmonics of degrees $k$ and $l$ respectively, then
$Y_{k}Y_{l}=\sum_{i=0}^{k+l}Z_{i},$
where each $Z_{i}$ is a spherical harmonic of degree $i$ $(i=0,1,\ldots,k+l)$.
Proof. We use polar coordinates $(r,\theta)$ on $\mathbb{R}^{n}$. Then $u_{k}(x)=r^{k}Q_{k}(\theta)$ and $u_{l}(x)=r^{l}Q_{l}(\theta)$ are homogeneous harmonic polynomials of degrees $k$ and $l$ respectively. Hence $u_{k}u_{l}$ is a homogeneous polynomial of degree $k+l$.
By the decomposition theorem for homogeneous polynomials stated in reference [5], we have
$u_{k}(x)u_{l}(x)=v_{k+l}(x)+|x|^{2}v_{k+l-2}(x)+\cdots+|x|^{\,k+l-\tau}v_{\tau}(x),$
where $\tau=1$ if $k+l$ is odd and $\tau=0$ if $k+l$ is even, and each $v_{i}$ is a homogeneous harmonic polynomial of degree $i$ $(i=k+l,\,k+l-2,\,\ldots,\,\tau)$. Restricting the above identity to the unit sphere yields Lemma 3.6.
[…]
$|\mathcal{L}z|=|f(z)|\leq Cz^{2},$
we now decompose the index set $\mathcal{I}$.
Define
$\mathcal{I}_{\rho}=\{\rho_{i}:i\geq 1\},$ (37)
and
$\mathcal{I}_{\widetilde{\rho}}=\left\{\sum_{i=1}^{r}n_{i}\rho_{i}:n_{i}\in\mathbb{Z}_{+},\;\sum_{i=1}^{r}n_{i}\geq 2\right\}.$ (38)
We assume that the sequence in $\mathcal{I}_{\widetilde{\rho}}$ is $\{\widetilde{\rho}_{i}\}_{i\geq 1}$, which is strictly increasing with $\widetilde{\rho}_{1}=2\rho_{1}$. We first consider the case where $\mathcal{I}_{\rho}\cap\mathcal{I}_{\widetilde{\rho}}$$\emptyset$; that is, no $\rho_{i}$ can be expressed as a positive-integer linear combination of $\rho_{1},\ldots,\rho_{i-1}$ (except for the trivial identity $\rho_{\widetilde{i}}=\rho_{i}$). In this situation, the elements of $\mathcal{I}$ may be arranged as follows:
$\rho_{1}\leq\cdots\leq\rho_{r_{1}}<\widetilde{\rho}_{1}<\cdots<\widetilde{\rho}_{l_{1}}<\rho_{r_{1}+1}\leq\cdots\leq\rho_{r_{2}}<\widetilde{\rho}_{l_{1}+1}<\cdots.$ (39)
For each $\widetilde{\rho}_{i}\in\mathcal{I}_{\widetilde{\rho}}$, we consider nonnegative integers $n_{1},\ldots,n_{r_{1}}$ such that
$n_{1}+\cdots+n_{r_{1}}\geq 2,\qquad n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}}=\widetilde{\rho}_{i}.$ (40)
Clearly, only finitely many such $(n_{1},\ldots,n_{r_{1}})$ exist.
Define
$\widetilde{R}_{i}=\max\Big\{$ $n_{1}\deg(Q_{1})+n_{2}\deg(Q_{2})+\cdots+n_{r_{1}}\deg(Q_{r_{1}}):$
[…]
and
$\widetilde{M}_{i}=\max\{\,m:\deg(Q_{m})\leq\widetilde{R}_{i}\,\}.$ (41)
We obtain From Lemma 2.3 that
$z=O\left(e^{\rho_{1}t}\right)$
Hence
$|\mathcal{L}(z)|=O(e^{2\rho_{1}t})=O(e^{\widetilde{\rho}_{1}t}).$
We now proceed in several steps to establish the case $\mathcal{I}_{\rho}\cap\mathcal{I}_{\widetilde{\rho}}$$\varnothing$.
Step 1. Observe that $\rho_{r_{1}}<\widetilde{\rho}_{1}=2\rho_{1}$. Thus, by Lemma A.8 in [4] (although the statement there is for $t\to+\infty$, the conclusion clearly remains valid as $t\to$-$\infty$; we will not repeat this remark later), there exists a function $\eta_{1}$ such that
$z=\eta_{1}+O(e^{\widetilde{\rho}_{1}t}),$
where
$\eta_{1}(t,\theta)=\sum_{i=1}^{r_{1}}c_{i}Q_{i}(\theta)e^{\rho_{i}t}.$ (42)
Define
$z_{1}=z-\eta_{1}.$ (43)
Then $\mathcal{L}\eta_{1}=0$, $\mathcal{L}z_{1}=f(z)$, and
$z_{1}=O(e^{\widetilde{\rho}_{1}t}).$ (44)
Step 2. We next show that there exists a function $\widetilde{\eta}_{1}$ such that
$\widetilde{z}_{1}=z_{1}-\widetilde{\eta}_{1}=z-\eta_{1}-\widetilde{\eta}_{1},$ (45)
and
$\mathcal{L}\widetilde{z}_{1}=O(e^{\widetilde{\rho}_{l_{1}+1}t}).$ (46)
We claim that $\widetilde{\eta}_{1}$ has the structure
$\widetilde{\eta}_{1}(t,\theta)=\sum_{i=1}^{l_{1}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}c_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t},$ (47)
where $\widetilde{M}_{i}$ is as defined in (41), and $c_{im}$ are constants. This provides the next level of expansion.
To prove this, fix $\widetilde{\eta}_{1}$ as above and define $\widetilde{z}_{1}$. Then
$\mathcal{L}\widetilde{z}_{1}=f(z)-\mathcal{L}\widetilde{\eta}_{1}.$ (48)
Note that $3\rho_{1}\in\mathcal{I}_{\widetilde{\rho}}$. We divide the discussion into two cases.
Case 1. Assume $\rho_{r_{1}+1}<3\rho_{1}$. Then $\widetilde{\rho}_{l_{1}}<\rho_{r_{1}+1}<3\rho_{1}$, which implies $\widetilde{\rho}_{l_{1}+1}\leq 3\rho_{1}$.
Observe that
$f(z)=f(z_{1}+\eta_{1})=\sum_{i=2}^{\infty}c_{i}(z_{1}+\eta_{1})^{i}.$
we get from (43) that
$z_{1}^{2}\leq Ce^{4\rho_{1}t},\qquad|z_{1}\eta_{1}|\leq Ce^{3\rho_{1}t}.$
From the expression of $\eta_{1}$ in (42), we obtain
$\sum_{i=2}^{\infty}c_{i}\eta_{1}^{i}=\sum_{n_{1}+\cdots+n_{r_{1}}\geq 2}a_{n_{1}\cdots n_{r_{1}}}e^{(n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}})t}Q_{1}^{n_{1}}\cdots Q_{r_{1}}^{n_{r_{1}}}.$
By the definition of $\mathcal{I}_{\widetilde{\rho}}$, it exists that $\widetilde{\rho}_{i}=n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}}$. Thus, by Lemma 3.6,
$\sum_{i=2}^{\infty}c_{i}\eta_{1}^{i}=\sum_{i=1}^{\infty}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}$
We now truncate the summation on the right-hand side at the finite index $l_{1}$ and denote the expression by $I_{1}$. Then
$I_{1}=\sum_{i=1}^{l_{1}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}.$
Hence,
$f(z)=I_{1}+O(e^{\widetilde{\rho}_{l_{1}+1}t}).$
Therefore, by (48),
$\mathcal{L}z_{1}=\mathcal{L}\widetilde{\eta}_{1}-I_{1}+O(e^{\widetilde{\rho}_{l_{1}+1}t}).$
We choose $\widetilde{\eta}_{1}$ to be the form
$\widetilde{\eta}_{1}(t,\theta)=\sum_{i=1}^{l_{1}}\sum_{m=0}^{\widetilde{M}_{i}}\widetilde{\eta}_{im}(t)Q_{m}(\theta).$
To solve the equation $\mathcal{L}\widetilde{\eta}_{1}=I_{1}$.we impose
$L_{m}\widetilde{\eta}_{im}=a_{im}e^{-\widetilde{\rho}_{i}t}.$ (49)
for each $1\leq i\leq l_{1}$ and $0\leq m\leq\widetilde{M}_{i}$.
Since $\rho_{m}\neq\widetilde{\rho}_{i}$ for every $m\neq i$, we can get from Lemma A.2 and Remark A.5 in [4](constants are viewed as special periodic functions) that
$\widetilde{\eta}_{im}(t)=c_{im}e^{\widetilde{\rho}_{i}t}.$ (50)
Thus we have obtained explicit formulas for $\widetilde{\eta}_{1}$ and $z_{1}$. Moreover, by (44) and (47), we obtain
$\widetilde{z}_{1}=O(e^{\widetilde{\rho}_{1}t}).$
Case 2. Assume now that $\rho_{r_{1}+1}>3\rho_{1}$. Then $\widetilde{\rho}_{l_{1}}\geq 3\rho_{1}$. Let $n_{1}$ be the largest integer satisfying $\widetilde{\rho}_{n_{1}}<3\rho_{1}$. Then $\widetilde{\rho}_{n_{1}+1}=3\rho_{1}$.
We repeat the argument of Case 1 with $n_{1}$ replacing $l_{1}$, and redefine $I_{1}$ so that the summation runs from $i=1$ to $n_{1}$. Similar to the expression for $\widetilde{\eta}_{1}$ in (47), we define
$\widetilde{\eta}_{11}(t,\theta)=\sum_{i=1}^{n_{1}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}c_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}.$ (51)
Set
$\widetilde{z}_{11}=z_{1}-\widetilde{\eta}_{11}.$
Then, by the same reasoning,
$\mathcal{L}\widetilde{z}_{11}=O(e^{\widetilde{\rho}_{n_{1}+1}t})=O(e^{3\rho_{1}t}).$
Combining this with (44) and (45), we obtain
$z_{11}=O\left(e^{\tilde{\rho}_{1}t}\right)=O\left(e^{2\rho_{1}t}\right)$
Since no $\rho_{i}$ lies between $\widetilde{\rho}_{1}$ and $\widetilde{\rho}_{r_{1}+1}$, Lemma A.8(ii) of [4] implies that
$z_{11}=O(e^{3\rho_{1}t}).$
We repeat a procedure similar to Step 2, replacing $\widetilde{\rho}_{1}=2\rho_{1}$ by $\widetilde{\rho}_{r_{1}+1}=3\rho_{1}$. If $\rho_{r_{1}+1}<4\rho_{1}$, we imitate the argument of Case 1; if $\rho_{r_{1}+1}>4\rho_{1}$, we imitate the argument of Case 2, selecting the largest integer $n_{2}$ such that $\widetilde{\rho}_{n_{2}}<4\rho_{1}$. Repeating this procedure a finite number of times and up to $\widetilde{\rho}_{l_{1}}$.
Step 3. As in Step 1, we now replace $\widetilde{\rho}_{1}$ with $\widetilde{\rho}_{l_{1}+1}$, and $1,\,r_{1},\,$ and $1$ with $r_{1}+1,\,r_{2},\,$ and $l_{1}+1$. Since $\rho_{r_{2}}<\widetilde{\rho}_{l_{1}+1}$, it follows from (46) and Lemma A.8(ii) of [4]that
$z_{1}(t,\theta)=\sum_{i=r_{1}+1}^{r_{2}}c_{i}Q_{i}(\theta)e^{\rho_{i}t}+O(e^{\widetilde{\rho}_{l_{1}+1}t}).$
we can discard the terms $e^{\rho_{i}t}$ for $i=1,\ldots,r_{1}$ since $\widetilde{z}_{1}=O(e^{\widetilde{\rho}_{1}t}).$ Define
$\eta_{2}(t,\theta)=\sum_{i=r_{1}+1}^{r_{2}}c_{i}X_{i}(\theta)e^{\rho_{i}t},$
and set
$z_{2}=\widetilde{z}_{1}-\eta_{2}.$
Then $\mathcal{L}\eta_{2}=0$ and
$z_{2}=z-\eta_{1}-\widetilde{\eta}_{1}-\eta_{2},\qquad z_{2}=O(e^{\widetilde{\rho}_{l_{1}+1}t}).$
Step 4. We proceed similarly to Step 2. Assume a function $\widetilde{\eta}_{2}$ is chosen and define
$\widetilde{z}_{2}=z_{2}-\widetilde{\eta}_{2}.$ (52)
Then
$\mathcal{L}\tilde{z}_{2}=f(z)-\mathcal{L}\tilde{\eta}_{1}-\mathcal{L}\tilde{\eta}_{2}$
Note that
$f(z)=f(z_{2}+\eta_{1}+\widetilde{\eta}_{1}+\eta_{2})=\sum_{i=2}^{\infty}c_{i}\left(z_{2}+\eta_{1}+\widetilde{\eta}_{1}+\eta_{2}\right)^{i}.$
We analyze $\sum_{i=2}^{\infty}c_{i}\left(\eta_{1}+\widetilde{\eta}_{1}+\eta_{2}\right)^{i}$ as in Step 2.Recall from Step 2 and (47) that by choosing $\widetilde{\eta}_{1}$ appropriately, we used $\mathcal{L}\widetilde{\eta}_{1}$ to cancel the terms $e^{\widetilde{\rho}_{i}t}$ in $f(z)$ for $i=1,\ldots,l_{1}$. Proceeding similarly, we can find a function $\widetilde{\eta}_{2}$ of the form
$\widetilde{\eta}_{2}(t,\theta)=\sum_{i=l_{1}+1}^{l_{2}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}c_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}$
to cancel the terms $e^{\widetilde{\rho}_{i}t}$ in $f(z)$ for $i=l_{1}+1,\ldots,l_{2}$.
As indicated in (52), defining $\widetilde{z}_{2}$ accordingly, we obtain
$\mathcal{L}z_{2}=O\left(e^{\widetilde{\rho}_{l_{2}+1}t}\right).$
Repeating the argument above completes the proof in the case $\mathcal{I}_{\rho}\cap\mathcal{I}_{\widetilde{\rho}}$$\emptyset$.
We now consider the more general situation in which some $\rho_{i}$ can be expressed as a positive-integer linear combination of $\rho_{1},\ldots,\rho_{i-1}$. Exponential terms in $t$ will appear in the solution of the equation $L_{i}\phi_{i}=a_{i}$ when a value $\rho_{i}$ coincides with a certain $\widetilde{\rho}_{i^{\prime}}$. According to Lemma A.2 of [4], such exponential items will appear one after another during the iteration process.
As a concrete example, we assume $\rho_{r_{1}}=\widetilde{\rho}_{1}$, replacing the strict inequality in (39). This is the first instance where some $\rho_{i}$ becomes equal to a $\widetilde{\rho}_{i^{\prime}}$.
In this case, we still have
$z=O(e^{\widetilde{\rho}_{1}t})$
and
$\mathcal{L}z=O(e^{2\rho_{1}t})=O(e^{\widetilde{\rho}_{1}t}).$
Following the same procedure as in Step 1, choose an index $r_{*}\in\{1,\ldots,r_{1}-1\}$ such that
$\rho_{r_{*}}<\rho_{r_{*}+1}=\cdots=\rho_{r_{1}}=\widetilde{\rho}_{1}=2\rho_{1}.$
By Lemma A.8(ii) of [4], we obtain
$z(t,\theta)=\sum_{i=1}^{r_{*}}c_{i}Q_{i}(\theta)e^{\rho_{i}t}+O\!\left(te^{\widetilde{\rho}_{1}t}\right).$
Define
$\eta_{1}(t,\theta)=\sum_{i=1}^{r_{*}}c_{i}X_{i}(\theta)e^{-\rho_{i}t}.$
With $z_{1}$ defined similarly to (43), we have
$z_{1}=O\!\left(te^{\widetilde{\rho}_{1}t}\right).$
Next, proceeding similarly to Step 2, for each $1\leq i\leq l_{1}$ and $0\leq m\leq\widetilde{M}_{i}$, we solve equation (49). If $\rho_{m}\neq\widetilde{\rho}_{i}$, then $\widetilde{\eta}_{im}(t)$ has the same expression as in (50). If $\rho_{m}=\widetilde{\rho}_{i}$, then $\widetilde{\eta}_{im}$ takes the form
$\widetilde{\eta}_{im}(t)=c_{i1m}\,t\,e^{\widetilde{\rho}_{i}t}+c_{i0m}\,e^{\widetilde{\rho}_{i}t}.$ (53)
Using the definition of $\widetilde{\eta}_{1}$ in (47), the new expression for $\widetilde{\eta}_{im}(t)$ in (53), and $\widetilde{z}_{1}$ in (45), we obtain (46). Repeating the same argument as above yields the desired result.
Now we discard multiple numbers and define a new index sequence $\{\mu_{i}\}_{i\geq 1}$. Clearly, $\mu_{1}=\rho_{1}=1$ and $\mu_{2}=\min\{2\rho_{1},\;\rho_{n+1}\}$.
We set
$\phi_{m}(t,\theta)=\sum_{\rho_{i}\leq\mu_{m}}c_{i}Q_{i}(\theta)e^{\rho_{i}t},$
and
$\widetilde{\phi}_{m}(t,\theta)=\sum_{\widetilde{\rho}_{i}\leq\mu_{m}}\sum_{j=0}^{i}\left\{\sum_{l=0}^{\widetilde{M}_{i}}c_{ijl}Q_{l}(\theta)\right\}t^{j}e^{\widetilde{\rho}_{i}t}.$
We note that $\phi_{m}$ is a solution of $\mathcal{L}\phi_{m}=0$ and $\widetilde{\phi}_{m}$ arises from the nonlinear term $f(z)$. In the special case that $\mathcal{I}_{\rho}\cap\mathcal{I}_{\widetilde{\rho}}$$\emptyset$,
$\widetilde{\phi}_{m}(t,\theta)=\sum_{\widetilde{\rho}_{i}\leq\mu_{m}}\left\{\sum_{l=0}^{\widetilde{M}_{i}}c_{il}Q_{l}(\theta)\right\}e^{\widetilde{\rho}_{i}t}.$
This completes the proof of (29) in Theorem 3.1. We now discuss the remaining cases. We will not present the full details, and for convenience of notation, the sequence $\{\rho_{i}\}_{i\geq 1}$ is again understood without multiple numbers. Accordingly, the index sets $\mathcal{I}$, $\mathcal{I}_{\rho}$, and $\mathcal{I}_{\widetilde{\rho}}$ can be defined in the same way (note that these indices differ from those used earlier).
The ordering of the sequences $\{\rho_{i}\}_{i\geq 1}$ and $\{\widetilde{\rho}_{i}\}_{i\geq 1}$ becomes
$\rho_{1}<\cdots<\rho_{r_{1}}\leq\widetilde{\rho}_{1}<\cdots<\widetilde{\rho}_{l_{1}}\leq\rho_{r_{1}+1}<\cdots<\rho_{r_{2}}\leq\widetilde{\rho}_{l_{1}+1}<\cdots.$ (54)
For case (30), we set
$\widetilde{M}_{i}=\max\{$ $kn_{1}+(k+1)n_{2}+\cdots+(k+i_{1}-1)n_{r_{1}}:$
[…]
For case (31), we set
[…]
For case (32), we set
[…]
For case (33), we set
[…]
For case (34), we set
$\widetilde{M}_{i}=\max\{$ $kn_{1}+(k+1)n_{2}+\cdots+jn_{j-k+1}+0\,n_{j-k+2}+(j+1)n_{j-k+3}+\cdots+(r_{1}+k-2)n_{r_{1}}:\;$
[…]
For case (35), we set
[…]
For case (36), we set
[…]
All remaining steps follow the same pattern as in the proof of (29). Thus Theorem 3.1 is now proved in full.
Remark. When $F\neq 0$, one may rewrite the equation in the form $\mathcal{L}\!\left(z+Ce^{\frac{2p-\alpha}{p+1}t}\right)=f(z)$ for a suitable constant $C$. Consequently, the expansion will contain terms involving $e^{\frac{2p-\alpha}{p+1}t}$.
4 Existence of Solutions
To establish the existence of solutions to (3), we introduce a weighted Hölder space. In this space, we apply the contraction mapping principle so that the fixed-point leads the existence of a solution to (3). Define
$\|v\|_{C_{\mu}^{i}\left((-\infty,t_{0}]\times S^{N-1}\right)}=\sum_{j=0}^{i}\sup_{(t,\theta)\in(-\infty,t_{0}]\times S^{N-1}}e^{-\mu t}\left|\nabla^{j}v(t,\theta)\right|,$
and
$\|v\|_{C_{\mu}^{i,a}\left((-\infty,t_{0}]\times S^{N-1}\right)}=\|v\|_{C_{\mu}^{i}\left((-\infty,t_{0}]\times S^{N-1}\right)}+\sup_{t\leq t_{0}-1}e^{-\mu t}\left[\nabla^{i}v\right]_{C^{a}\!\left([t-1,t+1]\times S^{N-1}\right)},$
where $\cdot]_{C^{a}}$ means the Hölder seminorm.
Definition. The weighted Hölder space $C_{\mu}^{i,\alpha}\left((-\infty,t_{0}]\times S^{N-1}\right)$ consists of those functions $v\in C^{i}\left((-\infty,t_{0}]\times S^{N-1}\right)$ for which the norm $\|v\|_{C_{\mu}^{i,\alpha}\left((-\infty,t_{0}]\times S^{N-1}\right)}$ is finite.
Let $\mathcal{L}$ be the linear operator defined in (7) and let $\mu>0$. For a function $g\in C_{\mu}^{0,\alpha}\left((-\infty,t_{0}]\times S^{N-1}\right)$, we consider the linear equation
$\mathcal{L}v=g.$ (55)
We shall impose a suitable boundary condition at $t=t_{0}$ so that
$\mathcal{L}:C_{\mu}^{2,\alpha}\!\left((-\infty,t_{0}]\times S^{N-1}\right)\;\longrightarrow\;C_{\mu}^{0,\alpha}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$
admits a bounded inverse. We begin with the Dirichlet boundary-value problem
$\begin{cases}\mathcal{L}v=g,&\text{in }(-\infty,t_{0})\times S^{N-1},\\[5.69054pt] v=\varphi,&\text{on }\{t_{0}\}\times S^{N-1}.\end{cases}$ (56)
Lemma 4.1.
Let $\mu>0$, $g\in C_{\mu}^{0}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$, and $\varphi\in C^{0}(S^{N-1})$. Then the above problem admits at most one solution $v\in C_{\mu}^{2}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$.
Proof. We give the argument for the case $\delta^{(k_{0}-1)}<\alpha<\delta^{(k_{0})}$. Let $g=0$ and $\varphi=0$, suppose $v\in C_{\mu}^{2}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$ is a solution of the problem. For each $k\geq 0$, define
$v_{k}(t)=\int_{S^{N-1}}v(t,\theta)\,Q_{k}(\theta)\,d\theta.$
Then $\mathcal{L}_{k}(v_{k})=0$ and $v_{k}(t_{0})=0$. Hence $v_{k}$ is a linear combination of elements in $\operatorname{Ker}(\mathcal{L}_{k})$. For $k<k_{0}$, we have
$v_{k}(t)=c_{k}^{1}e^{\Re(\sigma_{1}^{(k)})t}\cos(\gamma t)+c_{k}^{2}e^{\Re(\sigma_{1}^{(k)})t}\sin(\gamma t),$
or
$c_{k}^{1}e^{\sigma_{1}^{(k)}t}+c_{k}^{2}t\,e^{\sigma_{1}^{(k)}t},$
or
$c_{k}^{1}e^{\sigma_{1}^{(k)}t}+c_{k}^{2}e^{\sigma_{2}^{(k)}t},$
(where $\Re(\sigma_{1}^{(k)})<0$ and $\sigma_{1}^{(k)},\sigma_{2}^{(k)}\leq 0$).
From the assumption that $\lim_{t\to-\infty}v_{k}(t)=0$ and $v_{k}(t_{0})=0$, it follows immediately that $v_{k}(t)\equiv 0$.
When $k\geq k_{0}$, we have
$v_{k}(t)=c_{k}^{1}e^{\sigma_{1}^{(k)}t}+c_{k}^{2}e^{\sigma_{2}^{(k)}t}.$
Applying the Green identity, we obtain
$\int_{-\infty}^{t_{0}}\left[(\partial_{t}v_{k})^{2}-\frac{1}{2}\left(N-2+\frac{2\alpha+4}{p+1}\right)(v_{k}^{2})_{t}+\Big(\lambda_{k}-(\alpha+2)\big(N-2+\tfrac{\alpha+2}{p+1}\big)\Big)v_{k}^{2}\right]dt=0.$
Hence,
$\int_{-\infty}^{t_{0}}\left[(\partial_{t}v_{k})^{2}+\Big(\lambda_{k}-(\alpha+2)\big(N-2+\tfrac{\alpha+2}{p+1}\big)\Big)v_{k}^{2}\right]dt=0.$
For $k\geq k_{0}$, the coefficient
$\lambda_{k}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0.$
Thus $v_{k}(t)\equiv 0$. Therefore $v\equiv 0$.
We now estimate the $C^{2,\alpha}$ norm of the solution to (56).
Lemma 4.2.
Let $\alpha\in(0,1)$, $\mu>0$, $g\in C_{\mu}^{0,\alpha}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$ and $\varphi\in C^{2,\alpha}(S^{N-1})$. Suppose $v\in C_{\mu}^{2,\alpha}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$ is a solution of (56). Then
$\|v\|_{C_{\mu}^{2,\alpha}\left((-\infty,t_{0}]\times S^{N-1}\right)}\leq C\Big[\|v\|_{C_{\mu}^{0}\left((-\infty,t_{0}]\times S^{N-1}\right)}+\|g\|_{C_{\mu}^{0,\alpha}\left((-\infty,t_{0}]\times S^{N-1}\right)}+e^{-\mu t_{0}}\|\varphi\|_{C^{2,\alpha}(S^{N-1})}\Big],$ (57)
where $C>0$ depends only on $N,\alpha,\mu$ and is independent of $t_{0}$.
Proof. Fix an arbitrary $t\leq t_{0}$ and consider two cases.
(i) $t<t_{0}-2$. By the interior Schauder estimate,
$\sum_{j=0}^{2}\sup_{S^{N-1}}|\nabla^{j}v(t,\cdot)|+\left[\nabla^{2}v\right]_{C^{\alpha}([t-1,t+1]\times S^{N-1})}$
$\qquad\leq C\Big[\|v\|_{L^{\infty}([t-2,t+2]\times S^{N-1})}+\|g\|_{L^{\infty}([t-2,t+2]\times S^{N-1})}+[g]_{C^{\alpha}([t-2,t+2]\times S^{N-1})}\Big],$
where $C>0$ is independent of $t$.
To estimate the Hölder seminorm of $g$ on $t-2,t+2]\times S^{N-1}$, take arbitrary $(t_{1},\theta_{1}),(t_{2},\theta_{2})\in[t-2,t+2]\times S^{N-1}$ with $(t_{1},\theta_{1})\neq(t_{2},\theta_{2})$. Then we split into two cases: $|t_{1}-t_{2}|\leq 2$ and $|t_{1}-t_{2}|>2$. We present the proof of the first case.
If $|t_{1}-t_{2}|\leq 2$, then there exists some $t^{\prime}\in[t-1,t+1]$ such that $t_{1},t_{2}\in[t^{\prime}-1,t^{\prime}+1]$. Hence,
$g]_{C^{\alpha}([t-2,t+2]\times S^{N-1})}\leq\max\Big\{\sup_{t^{\prime}\in[t-1,t+1]}[g]_{C^{\alpha}([t^{\prime}-1,t^{\prime}+1]\times S^{N-1})},\ \|g\|_{L^{\infty}([t-2,t+2]\times S^{N-1})}\Big\}.$
Thus,
$\sum_{j=0}^{2}\sup_{S^{N-1}}\left|\nabla^{j}v(t,\cdot)\right|+\left[\nabla^{2}v\right]_{C^{\alpha}\left([t-1,t+1]\times S^{N-1}\right)}$
$\quad\leq C\left[\|v\|_{L^{\infty}\left([t-2,t+2]\times S^{N-1}\right)}+\|g\|_{L^{\infty}\left([t-2,t+2]\times S^{N-1}\right)}+\sup_{t^{\prime}\in[t-1,t+1]}[g]_{C^{\alpha}\left(\left[t^{\prime}-1,t^{\prime}+1\right]\times S^{N-1}\right)}\right]$
Multiplying both sides by $e^{-\mu t}$ and then taking the supremum over $t\in(-\infty,t_{0}-2)$, we obtain
$\sum_{j=0}^{2}\sup_{t\in(-\infty,t_{0}-2)}\sup_{S^{N-1}}e^{-\mu t}\left|\nabla^{j}v(t,\cdot)\right|+\sup_{t\in(-\infty,t_{0}-2)}e^{-\mu t}\left[\nabla^{2}v\right]_{C^{\alpha}([t-1,t+1]\times S^{N-1})}$
$\qquad\leq C\left[\|v\|_{C_{\mu}^{0}((-\infty,t_{0}]\times S^{N-1})}+\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})}\right],$
where the constant $C>0$ is independent of $t_{0}$.
(ii) $t_{0}-2\leq t\leq t_{0}$. By the boundary Schauder estimate,
$\sum_{j=0}^{2}\sup_{S^{N-1}}\left|\nabla^{j}v(t,\cdot)\right|+\left[\nabla^{2}v\right]_{C^{\alpha}([t_{0}-3,t_{0}]\times S^{N-1})}$
$\qquad\leq C\left[\|v\|_{L^{\infty}([t_{0}-4,t_{0}]\times S^{N-1})}+\|g\|_{L^{\infty}([t_{0}-4,t_{0}]\times S^{N-1})}+[g]_{C^{\alpha}([t_{0}-4,t_{0}]\times S^{N-1})}+\|\varphi\|_{C^{2,\alpha}(S^{N-1})}\right].$
Using an argument similar to the one above, we obtain
$\sum_{j=0}^{2}\sup_{t\in[t_{0}-2,t_{0}]}\sup_{S^{N-1}}e^{-\mu t}\left|\nabla^{j}v(t,\cdot)\right|+\sup_{t\in[t_{0}-2,t_{0}-1]}e^{-\mu t}\left[\nabla^{2}v\right]_{C^{\alpha}([t-1,t+1]\times S^{N-1})}$
$\qquad\leq C\left[\|v\|_{C_{\mu}^{0}((-\infty,t_{0}]\times S^{N-1})}+\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})}+e^{-\mu t_{0}}\|\varphi\|_{C^{2,\alpha}(S^{N-1})}\right].$
Combining the two cases above yields the desired estimate
Lemma 4.3.
Assume $\mu>0$ and $\mu\neq\sigma_{1}^{(k)}$ for every $k\geq 1$. Let $T$ and $t_{0}$ be constants with $t_{0}\leq 0$ and $T-t_{0}\leq-4$, suppose $g\in C^{0}([T,t_{0}]\times S^{N-1})$. $v\in C^{2}([T,t_{0}]\times S^{N-1})$ satisfies
$\begin{cases}\mathcal{L}v=g&\text{in }(T,t_{0})\times S^{N-1},\\ v=0&\text{on }(\{T\}\cup\{t_{0}\})\times S^{N-1},\end{cases}$
and for every $t\in[T,t_{0}]$ and every $k=0,1,2,\ldots,K$
$\int_{S^{N-1}}v(t,\theta)\,Q_{k}(\theta)\,d\theta=0,$
where $K$ is the largest integer such that $\sigma_{1}^{(K)}<\mu$, then
$\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}|v(t,\theta)|\leq C\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}|g(t,\theta)|,$
where $C>0$ depends only on $N$ and $\mu$ and independent of $T$ and $t_{0}$.
Proof. Assume the conclusion is false. Then there exist sequences $\{T_{i}\}$, $\{t_{i}\}$, $\{v_{i}\}$ and $\{g_{i}\}$ with $t_{i}\leq 0$ and $T_{i}-t_{i}\leq-4$, such that
$\begin{cases}\mathcal{L}v_{i}=g_{i}&\text{in }(T_{i},t_{i})\times S^{N-1},\\ v_{i}=0&\text{on }(\{T_{i}\}\cup\{t_{i}\})\times S^{N-1},\end{cases}$
and
$\sup_{(t,\theta)\in[T_{i},t_{i}]\times S^{N-1}}e^{-\mu t}|g_{i}(t,\theta)|=1,\qquad\sup_{(t,\theta)\in[T_{i},t_{i}]\times S^{N-1}}e^{-\mu t}|v_{i}(t,\theta)|\to\infty.\quad(\text{as }i\to\infty.)$
From each interval $(T_{i},t_{i})$ ,choose a point $t_{i}^{*}\in(T_{i},t_{i})$ such that
$M_{i}=\sup_{S^{N-1}}e^{-\mu t_{i}^{*}}\left|v_{i}(t_{i}^{*},\cdot)\right|=\sup_{(t,\theta)\in[T_{i},t_{i}]\times S^{N-1}}e^{-\mu t}|v_{i}(t,\theta)|.$
Then $M_{i}\to\infty$ as $i\to\infty$. Define
$\widetilde{v}_{i}(t,\theta)=M_{i}^{-1}e^{-\mu t_{i}^{*}}v_{i}(t+t_{i}^{*},\theta),$
and
$\widetilde{g}_{i}(t,\theta)=M_{i}^{-1}e^{-\mu t_{i}^{*}}g_{i}(t+t_{i}^{*},\theta).$
Then
$\sup_{S^{N-1}}|\widetilde{v}_{i}(0,\cdot)|=1,$
For every $(t,\theta)\in[T_{i}-t_{i}^{*},\,t_{i}-t_{i}^{*}]\times S^{N-1}$ we have
$\left|e^{-\mu t}\tilde{v}_{i}(t,\theta)\right|\leq 1.$ (58)
Furthermore, for all $t\in[T_{i}-t_{i}^{*},\,t_{i}-t_{i}^{*}]\times S^{N-1}$, we have
$\mathcal{L}\widetilde{v}_{i}=\widetilde{g}_{i}.$
Passing to a subsequence, we may assume that there exist $\tau_{-}\in\mathbb{R}^{-}\cup\{-\infty\}$ and $\tau_{+}\in\mathbb{R}^{+}\cup\{\infty\}$ such that
$T_{i}-t_{i}^{*}\to\tau_{-},\qquad t_{i}-t_{i}^{*}\to\tau_{+}.$
In fact, we obtain from (58) that
$|\widetilde{v}_{i}|\leq Ce^{\mu(t_{i}^{*}-T_{i})}\qquad\text{on }(T_{i}-t_{i}^{*},\,T_{i}-t_{i}^{*}+2)\times S^{N-1},$
hence
$\left|\frac{d^{2}\widetilde{v}_{i}}{dt^{2}}+\left(N-2+\frac{4}{p+1}\right)\frac{d\widetilde{v}_{i}}{dt}+\Delta_{\theta}\widetilde{v}_{i}\right|\leq Ce^{\mu(t_{i}^{*}-T_{i})}\qquad\text{on }(T_{i}-t_{i}^{*},\,T_{i}-t_{i}^{*}+2)\times S^{N-1}.$
Since $\widetilde{v}_{i}=0$ on $\{T_{i}-t_{i}^{*}\}\times S^{N-1}$, it follows that
$|\nabla\widetilde{v}_{i}|\leq Ce^{\mu(t_{i}^{*}-T_{i})}\qquad\text{on }(T_{i}-t_{i}^{*},\,T_{i}-t_{i}^{*}+1)\times S^{N-1}.$
Consequently $T_{i}-t_{i}^{*}$ keeps a definite distance away from 0. Similarly, $t_{i}-t_{i}^{*}$ also keeps a definite distance away from 0. Hence,
$0\in(\tau_{-},\tau_{+}).$
Assume
$\widetilde{v}_{i}\to\widehat{v}\quad\text{uniformly on compact subsets of }(\tau_{-},\tau_{+})\times S^{N-1}.$
Also $\widetilde{g}_{i}$ converges uniformly to 0 on every compact subset of $(\tau_{-},\tau_{+})\times S^{N-1}$. Therefore, $\widehat{v}\neq 0$, and
$|e^{-\mu t}\widehat{v}(t,\theta)|\leq 1,$ (59)
for all $(t,\theta)\in(\tau_{-},\tau_{+})\times S^{N-1}$ .Also we have
$\mathcal{L}\widehat{v}=0.$
on $(\tau_{-},\tau_{+})\times S^{N-1}$
Furthermore,
$\lim\widehat{v}(t,\theta)=0.$ (60)
Now, for any $k\geq 0$, define
$\widehat{v}_{k}(t)=\int_{S^{N-1}}\widehat{v}(t,\theta)\,Q_{k}(\theta)\,d\theta.$
Then $\mathcal{L}_{k}(\widehat{v}_{k})=0$, $\widehat{v}_{k}$ is a linear combination of a basis of $\operatorname{Ker}(\mathcal{L}_{k})$. Choose $k$ such that $\sigma_{1}^{(k)}>\mu>0$. Then
$\widehat{v}_{k}(t)=c_{k}^{1}e^{\sigma_{1}^{(k)}t}+c_{k}^{2}e^{\sigma_{2}^{(k)}t}.$
From (59) we know
$|e^{-\mu t}\widehat{v}_{k}(t)|\leq C.$
for all $t\in(\tau_{-},\tau_{+})$.
If $\tau_{+}=+\infty$, then necessarily $c_{k}^{1}=0$, thus $\widehat{v}_{k}(t)=c_{k}^{2}e^{\sigma_{2}^{(k)}t}$ which decays exponentially as $t\to+\infty$. If $\tau_{+}$ is finite,then $\lim_{t\to\tau_{+}}\widehat{v}_{k}(t)=0$ by (60),
Similarly, if $\tau_{-}=$-$\infty$, then $c_{k}^{2}=0$, hence $\widehat{v}_{k}(t)=c_{k}^{1}e^{\sigma_{1}^{(k)}t}$ which decays exponentially as $t\to$-$\infty$. If $\tau_{-}$ is finite, then $\lim_{t\to\tau_{-}}\widehat{v}_{k}(t)=0$ by (60),
Therefore,
$\int_{\tau_{-}}^{\tau_{+}}\left[(\partial_{t}\widehat{v}_{k})^{2}+\left(\lambda_{k}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)\right)\widehat{v}_{k}^{2}\right]dt=0.$
we have
$\lambda_{k}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0,$
Since $\sigma_{1}^{(k)}>\mu>0$. So $\widehat{v}_{k}\equiv 0$ for every $k$ satisfying $\sigma_{1}^{(k)}>\mu$ .
From the assumption that
$\int_{S^{N-1}}\widetilde{v}_{i}(t,\theta)Q_{k}(\theta)\,d\theta=0\qquad\text{for all }k=0,1,\ldots,K\text{ with }\sigma_{1}^{(K)}<\mu,$
since $\widetilde{v}_{i}\to\widehat{v}$ as $i\to\infty$, we deduce that
$\widehat{v}_{k}\equiv 0\qquad\text{for all }k=0,1,\ldots,K.$
Combining the above, we conclude that $\widehat{v}_{k}\equiv 0$ for all $k\geq 0$, therefore $\widehat{v}\equiv 0$, contradicting our earlier conclusion that $\widehat{v}\neq 0$. Lemma 4.3 is proved.
Lemma 4.4.
Let $\alpha\in(0,1)$ and $\mu>\sigma_{1}^{(K)}$ (with $K\geq 0$), and let $g\in C_{\mu}^{0,\alpha}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$ satisfying that $g(t,\cdot)\in\operatorname{span}\{Q_{0},Q_{1},\ldots,Q_{K}\}$ for all $t\leq t_{0}$. Then equation (55) admits a unique solution $v\in C_{\mu}^{2,\alpha}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$, and for every $t\leq t_{0}$, $v(t,\cdot)\in\operatorname{span}\{Q_{0},Q_{1},\ldots,Q_{K}\}.$ Moreover, the map $g\mapsto v$ is linear, and
$\|v\|_{C_{\mu}^{2,\alpha}\left((-\infty,t_{0}]\times S^{N-1}\right)}\leq C\|g\|_{C_{\mu}^{0,\alpha}\left((-\infty,t_{0}]\times S^{N-1}\right)},$
where the constant $C>0$ depends only on $N,\alpha,\mu$ and is independent of $t_{0}$.
Proof. For each $k=0,1,\ldots,K$, define
$g_{k}(t)=\int_{S^{N-1}}g(t,\theta)\,Q_{k}(\theta)\,d\theta.$
Then
$\|g_{k}\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}])}\leq C\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})},$
and
$g(t,\theta)=\sum_{k=0}^{K}g_{k}(t)\,Q_{k}(\theta).$
Let $\mathcal{L}_{k}$ denote the linear operator from (8). Consider the ordinary differential equation
$\mathcal{L}_{k}v_{k}=g_{k}.$ (61)
We claim that (61) admits a solution $v_{k}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}])$ satisfying
[…]
where the constant $C$ depends only on $N,\alpha,\mu$ and not on $t_{0}$. set
$v_{k}(t)=B_{k}^{1}\int_{-\infty}^{t}e^{\sigma_{1}^{(k)}(t-s)}g_{k}(s)ds-B_{k}^{2}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{k}(s)ds,$ (63)
or
$v_{k}(t)$ $=B_{k}^{1}\cos\bigl(I(\sigma_{1}^{(k)}t)\bigr)\int_{-\infty}^{t}e^{\sigma_{1}^{(k)}(t-s)}g_{k}(s)\sin\bigl(I(\sigma_{1}^{(k)}s)\bigr)\,ds$ (62)
$\quad-B_{k}^{2}\cos\bigl(I(\sigma_{1}^{(k)}t)\bigr)\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{k}(s)\sin\bigl(I(\sigma_{1}^{(k)}s)\bigr)\,ds,$
$\left|B_{k}^{1}\right|=\left|\frac{1}{\sigma_{2}^{(k)}-\sigma_{1}^{(k)}}\right|.$We now write down the argument for the case of (61).
A direct computation shows that, for all $t\leq t_{0}$,
$\begin{gathered}e^{-\mu t}|v_{k}(t)|\leq C\sup_{t\leq t_{0}}e^{-\mu t}|g_{k}(t)|=C\|g_{k}\|_{C_{\mu}^{0}((-\infty,t_{0}])},\\[5.69054pt] e^{-\mu t}\left(|v_{k}^{\prime}(t)|+|v_{k}^{\prime\prime}(t)|\right)\leq C\|g_{k}\|_{C_{\mu}^{0}((-\infty,t_{0}])}.\end{gathered}$
To estimate the Hölder seminorm of $v_{k}^{\prime\prime}$, decompose
$v_{k}^{\prime\prime}(t)=R_{1}(t)+R_{2}(t),$
where
$R_{1}(t)=B_{k}^{1}(e^{\sigma_{1}^{(k)}t})^{\prime\prime}\int_{-\infty}^{t}e^{-\sigma_{1}^{(k)}s}g_{k}(s)\,ds-B_{k}^{1}(e^{\sigma_{2}^{(k)}t})^{\prime\prime}\int_{-\infty}^{t}e^{-\sigma_{2}^{(k)}s}g_{k}(s)\,ds,$
and
$R_{2}(t)=B_{k}^{1}(e^{\sigma_{1}^{(k)}t})^{\prime}e^{-\sigma_{1}^{(k)}t}g_{k}(t)-B_{k}^{1}(e^{\sigma_{2}^{(k)}t})^{\prime}e^{-\sigma_{2}^{(k)}t}g_{k}(t).$
Thus,
$R_{1}^{\prime}(t)={}$ $B_{k}^{1}(e^{\sigma_{1}^{(k)}t})^{\prime\prime\prime}\int_{-\infty}^{t}e^{-\sigma_{1}^{(k)}s}g_{k}(s)\,ds-B_{k}^{1}(e^{\sigma_{2}^{(k)}t})^{\prime\prime\prime}\int_{-\infty}^{t}e^{-\sigma_{2}^{(k)}s}g_{k}(s)\,ds$
$+B_{k}^{1}(e^{\sigma_{1}^{(k)}t})^{\prime\prime}e^{-\sigma_{1}^{(k)}t}g_{k}(t)-B_{k}^{1}(e^{\sigma_{2}^{(k)}t})^{\prime\prime}e^{-\sigma_{2}^{(k)}t}g_{k}(t).$
Similarly, for all $t\leq t_{0}$,
$e^{-\mu t}|R_{1}^{\prime}(t)|\leq C\|g_{k}\|_{C_{\mu}^{0}((-\infty,t_{0}])},$
hence, for all $t\leq t_{0}-1$,
$e^{-\mu t}[R_{1}]_{C^{\alpha}([t-1,t+1])}$ $\leq C\|g_{k}\|_{C_{\mu}^{0}((-\infty,t_{0}])},$
$e^{-\mu t}[R_{2}]_{C^{\alpha}([t-1,t+1])}$ $\leq C\|g_{k}\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}])}.$
Therefore, for all $t\leq t_{0}-1$,
$e^{-\mu t}\left[v_{k}^{\prime\prime}\right]_{C^{\alpha}([t-1,t+1])}\leq C\|g_{k}\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}])}.$
Combining the above estimates yields (62).
After get the solution of $v_{k}$ from (61) for each $k=0,1,\ldots,K$, we set
$v(t,\theta)=\sum_{k=0}^{K}v_{k}(t)Q_{k}(\theta).$
Clearly $\mathcal{L}v=g$, then we can get from (62) that
$\|v\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}$ $\leq C\sum_{k=0}^{K}\|v_{k}\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}])}$
$\leq C\sum_{k=0}^{K}\|g_{k}\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}])}$
$\leq C\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})}.$
Thus $v$ is the desired solution. Moreover, it’s clear that such a solution is unique under the additional condition that $v(t,\cdot)\in\operatorname{span}\{Q_{0},Q_{1},\ldots,Q_{K}\}\text{for every }t\leq t_{0}.$
Lemma 4.5.
Let $\alpha\in(0,1)$, $\mu>0$ with $\mu\neq\sigma_{1}^{(k)}$ for every $k\geq 1$, and let $g\in C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})$ satisfying
$\int_{S^{N-1}}g(t,\theta)\,Q_{k}(\theta)\,d\theta=0,\qquad\forall\,t\leq t_{0},\quad k=0,1,\ldots,K,$
where $K$ is the largest integer such that $\sigma_{1}^{(K)}<\mu$. Then equation (55) admits a unique solution $v\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ satisfying $v=0$ on $\{t_{0}\}\times S^{N-1}$. Furthermore,
$\|v\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}\leq C\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})},$
where $C>0$ depends only on $N,\alpha,\mu$ and is independent of $t_{0}$.
Proof. Fix any $T\leq t_{0}-4$. We first show that there exists $v_{T}\in C^{2,\alpha}([T,t_{0}]\times S^{N-1})$ such that
$\begin{cases}\mathcal{L}v_{T}=g&\text{in }(T,t_{0})\times S^{N-1},\\[2.84526pt] v_{T}=0&\text{on }(\{T\}\cup\{t_{0}\})\times S^{N-1}.\end{cases}$ (64)
Observe that the problem (64) is equivalent to
$\begin{cases}\frac{\partial}{\partial t}\left(e^{\hat{A}t}\frac{\partial v_{T}}{\partial t}\right)+e^{\hat{A}t}\Delta_{\theta}v_{T}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)e^{\hat{A}t}v_{T}=e^{\hat{A}t}g&\text{ in }\left(T,t_{0}\right)\times S^{N-1},\\ v_{T}=0&\text{ on }\left(\{T\}\cup\{t_{0}\}\right)\times S^{N-1},\end{cases}$
Here $\hat{A}=N-2+\frac{2\alpha+4}{p+1}$. Consider the energy functional
$\mathcal{G}_{T}(v)=\int_{T}^{t_{0}}\int_{S^{N-1}}\left[e^{\hat{A}t}(\partial_{t}v)^{2}+e^{\hat{A}t}|\nabla_{\theta}v|^{2}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)e^{\hat{A}t}v^{2}+2e^{\hat{A}t}gv\right]dt\,d\theta.$
Define
$\Gamma=\left\{\phi\in H^{1}(S^{N-1}):\int_{S^{N-1}}\phi(\theta)\,Q_{k}(\theta)\,d\theta=0,\quad k=0,1,2,\ldots,K,\ \sigma_{1}^{(k)}<\mu\right\}.$
Since
$\int_{S^{N-1}}|\nabla_{\theta}\phi|^{2}\,d\theta\geq K(N+K-2)\int_{S^{N-1}}\phi^{2}\,d\theta,$
it follows that for any $v\in H_{0}^{1}((T,t_{0})\times S^{N-1})$ satisfying $v(t,\cdot)\in\Gamma$ ( $t\in(T,t_{0})$), we have
$\mathcal{G}_{T}(v)\geq\int_{T}^{t_{0}}\int_{S^{N-1}}e^{\hat{A}t}(\partial_{t}v)^{2}+\left(K(N+K-2)-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)\right)e^{\hat{A}t}v^{2}+2e^{\hat{A}t}gv~~dtd\theta.$
Since $\sigma_{1}^{(K)}>0$, we have
$K(N+K-2)-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0.$
Hence, the functional $\mathcal{G}_{T}$ is coercive and weakly lower semicontinuous. Therefore we may find a minimizer $v_{T}$ of $\mathcal{G}_{T}$ in the space
$\left\{v\in H_{0}^{1}((T,t_{0})\times S^{N-1}):v(t,\cdot)\in\Gamma\text{ for every }t\in(T,t_{0})\right\}.$
Since $g(t,\cdot)\in\Gamma$ for all $t\in(T,t_{0})$, it follows that $v_{T}$ is a solution of (64), and $v_{T}(t,\cdot)\in\Gamma$ for all $t\in(T,t_{0})$.
By Lemma 4.3, we have
$\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}\,|v_{T}(t,\theta)|\leq C\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}\,|g(t,\theta)|,$
where the constant $C>0$ depends only on $N$ and $\mu$ and is independent of $T$ and $t_{0}$.
For any fixed $T_{0}<t_{0}$, consider the region $t_{0}+T_{0},\,t_{0}]\times S^{N-1}\\ \subset[t_{0}+T_{0}-1,\,t_{0}]\times S^{N-1}$. By the interior and boundary Schauder estimates, together with the fact that $v_{T}(t_{0},\theta)=0$, we may extract a subsequence $v_{T}$ converges to a $C^{2,\alpha}$ solution $v$ of (55) on $t_{0}+T_{0},\,t_{0}]\times S^{N-1}$ (via the Arzelà–Ascoli theorem), with $v=0$ on $\{t_{0}\}\times S^{N-1}$ as $T\to$-$\infty$. Via a diagonalization process, we conclude that $v_{T}$ converges to a $C^{2,\alpha}$ solution $v$ of (55)on $(-\infty,t_{0}]\times S^{N-1}$ , with $v=0$ on $\{t_{0}\}\times S^{N-1}$.
Furthermore,
$\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}|v(t,\theta)|\leq C\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}|g(t,\theta)|,$
or
$\|v\|_{C_{\mu}^{0}((-\infty,t_{0}]\times S^{N-1})}\leq C\|g\|_{C_{\mu}^{0}((-\infty,t_{0}]\times S^{N-1})},$ (65)
where $C>0$ depends only on $N$ and $\mu$, and not on $t_{0}$.
Substituting (65) into (57) (with $\varphi=0$) and then completes the proof.
Theorem 4.6.
Let $\alpha\in(0,1)$ and $\mu>0$ with $\mu\neq\sigma_{1}^{(k)}$ for all $k\geq 1$, let $g\in C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})$. Then equation (55) admits a solution $v\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ satisfying
$\|v\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}\leq C\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})},$
where the constant $C>0$ depends only on $N,\alpha,\mu$ and is independent of $t_{0}$. Moreover, the corresponding map $g\mapsto v$ is linear.
Proof. Let $K\geq 0$ be the largest integer such that $\sigma_{1}^{(K)}<\mu$. For $k=0,1,\ldots,K$, define
$g_{k}(t)=\int_{S^{N-1}}g(t,\theta)\,Q_{k}(\theta)\,d\theta.$
Let $v_{1}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ be the unique solution (by Lemma 4.4) of
$\mathcal{L}(v_{1})=\sum_{k=0}^{K}g_{k}(t)\,Q_{k}(\theta)\qquad\text{in }(-\infty,t_{0}]\times S^{N-1}.$
Next, by Lemma 3.6, choose $v_{2}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ to be the unique solution of
$\begin{cases}\mathcal{L}v=g-\sum_{k=0}^{K}g_{k}Q_{k}&\text{in }(-\infty,t_{0}]\times S^{N-1},\\[5.69054pt] v=0&\text{on }\{t_{0}\}\times S^{N-1}.\end{cases}$
Then $v=v_{1}+v_{2}$ is the desired solution. Note that
$v(t_{0},\theta)=v_{1}(t_{0},\theta)=\sum_{k=0}^{K}v_{k}(t_{0})\,Q_{k}(\theta),$
where each $v_{k}(t)$ ($k=0,1,\ldots,K$) is obtained from Lemma 4.4.
Remark 4.7.
We denote by $\mathcal{L}^{-1}$ the correspondence $g\mapsto v$ in Theorem 4.6. Then
$\mathcal{L}^{-1}:C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})\longrightarrow C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$
is a bounded linear operator, and its operator norm is independent of $t_{0}$.
Next we prove the existence of solutions to (3). Let
$H(z)=z_{tt}+\left(N-2+2q\right)z_{t}+(\alpha+2)(N-2+q)\,z+\Delta_{\theta}z-f(z)-Fe^{\frac{2p-\alpha}{p+1}t}.$
It’s equivalent to prove that the equation $H(z)=0$ admits a solution.
Theorem 4.8.
(1) Radial case. Assume $\mu>0$ and $F\neq 0$. Suppose $\hat{z}\in C^{2,\alpha}((-\infty,0])$ satisfying
$|\hat{z}(t)|+|\hat{z}^{\prime}(t)|\to 0\quad\text{as }t\to-\infty,$
and there exists a positive constant $C$ such that
$\big|\mathcal{N}(\hat{z})\big|+\left|\frac{d}{dt}\mathcal{N}(\hat{z})\right|\leq Ce^{\mu t}\quad\forall\,t\in(-\infty,0].$
Then there exists $t_{0}<0$ and a solution $z(t)\in C^{2,\alpha}((-\infty,t_{0}])$ of H(z)=0{H}(z)=0 such that
[…]
where $C$ is a positive constant.
(2) Nonradial case. Assume $\mu>0$ and $\mu\neq\sigma_{1}^{(i)}$ for every $i\geq 1$. Suppose $\hat{z}\in C^{2,\alpha}((-\infty,0]\times\mathbb{S}^{N-1})$ satisfies
$|\hat{z}(t,\theta)|+|\nabla\hat{z}(t,\theta)|\to 0\quad\text{as }t\to-\infty\text{ uniformly for }\theta\in\mathbb{S}^{N-1},$
and there exists a positive constant $C$ such that for all $(t,\theta)\in(-\infty,0]\times\mathbb{S}^{N-1}$,
$|\mathcal{N}(\hat{z})|+|\nabla\mathcal{N}(\hat{z})|\leq Ce^{\mu t}.$ (66)
Then there exist $t_{0}<0$ and a solution
$z(t,\theta)\in C^{2,\alpha}\big((-\infty,t_{0}]\times\mathbb{S}^{N-1}\big)\quad\text{of }{H}(z)=0,$
such that
$|z(t,\theta)-\hat{z}(t,\theta)|\leq Ce^{\mu t}\quad\text{for }(t,\theta)\in(-\infty,t_{0})\times\mathbb{S}^{N-1},$
where $C$ is a positive constant.
Proof. We only prove case (2), since the radial case (1) is similar and easier. For any $\phi\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$, we have
$\mathcal{N}(\widehat{z}+\phi)=\mathcal{N}(\widehat{z})+\mathcal{L}\phi-P(\phi),$
where
$P(\phi)=(\widehat{z}+\Lambda+\phi)^{-p}-(\widehat{z}+\Lambda)^{-p}+p\Lambda^{-(p+1)}\phi.$
Thus $\mathcal{N}(\widehat{z}+\phi)=0$ is equivalent to
$\mathcal{L}\phi=-\big[\mathcal{N}(\widehat{z})-P(\phi)\big].$ (67)
By Theorem 3.7 and Remark 3.8, we may rewrite (66) in the form
$\phi=\mathcal{L}^{-1}\big(-\mathcal{N}(\widehat{z})+P(\phi)\big).$
Define the operator
$\mathcal{T}(\phi)=\mathcal{L}^{-1}\big(-\mathcal{N}(\widehat{z})+P(\phi)\big).$
We shall show that, for $t_{0}<0$ with $|t_{0}|$ sufficiently large, $\mathcal{T}$ is a contraction on a suitable ball in $C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$. Set
$\Gamma_{B,t_{0}}=\big\{z\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1}):\|z\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}\leq B\big\}.$
[…]
First, by (66) we have
$\|\mathcal{N}(\widehat{z})\|_{C_{\mu}^{1}((-\infty,t_{0}]\times S^{N-1})}\leq C_{1}.$
Define
$E(\phi)=(-p)\int_{0}^{1}\big[(\widehat{z}+\Lambda+s\phi)^{-(p+1)}-\Lambda^{-(p+1)}\big]\,ds.$ (68)
We get that $P(\phi)=\phi\,E(\phi)$.
Take any $\phi\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1}))$ satisfying $\|\phi\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}\leq B$. Note that
$|\widehat{z}|+|\nabla\widehat{z}|\leq\varepsilon(t),$
where $\varepsilon(t)$ is a monotonically increasing function with $\varepsilon(t)\to 0$ as $t\to$-$\infty$.
$|\phi|+|\nabla\phi|\leq Be^{\mu t}.$
Hence, for all $t\leq t_{0}$,
$|E(\phi)|+|\nabla E(\phi)|\leq C_{2}\big(\varepsilon(t)+Be^{\mu t}\big),$ (69)
therefore
$\|P(\phi)\|_{C_{\mu}^{1}((-\infty,t_{0}]\times S^{N-1})}\leq C_{2}\big(\varepsilon(t_{0})+Be^{\mu t_{0}}\big)\|\phi\|_{C_{\mu}^{1}((-\infty,t_{0}]\times S^{N-1})}\leq C_{2}\big(\varepsilon(t_{0})+Be^{\mu t_{0}}\big)\,B.$
By Theorem 3.7 we obtain
[…]
$\leq C\left[C_{1}+C_{2}\left(\epsilon\left(t_{0}\right)+Be^{\mu t_{0}}\right)B\right],$
$C,C_{1},C_{2}$ are all positive and independent of $t_{0}$. Choose $B\geq 2CC_{1}$ and $|t_{0}|$ sufficiently large so that
$CC_{2}\big(\varepsilon(t_{0})+Be^{\mu t_{0}}\big)\leq\frac{1}{2}.$
It follows that
$\|\mathcal{T}(\phi)\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}\leq B.$
This gives the required self-mapping property. We now show that $\mathcal{T}:\Gamma_{B,t_{0}}\to\Gamma_{B,t_{0}}$ is a contraction map; namely, for any $\phi_{1},\phi_{2}\in\Gamma_{B,t_{0}}$, there exists $\kappa\in(0,1)$ such that
$\|\mathcal{T}(\phi_{1})-\mathcal{T}(\phi_{2})\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}\leq\kappa\|\phi_{1}-\phi_{2}\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}.$ (70)
Observe that
$\mathcal{T}(\phi_{1})-\mathcal{T}(\phi_{2})=\mathcal{L}^{-1}\big(P(\phi_{1})-P(\phi_{2})\big),$
and
$P(\phi_{1})-P(\phi_{2})$ $=\phi_{1}E(\phi_{1})-\phi_{2}E(\phi_{2})$
$=(\phi_{1}-\phi_{2})E(\phi_{1})\,+\,\phi_{2}\big(E(\phi_{1})-E(\phi_{2})\big).$
By (68),
$E(\phi_{1})-E(\phi_{2})=(-p)\int_{0}^{1}\big[(\widehat{z}+\Lambda+s\phi_{1})^{-(p+1)}-(\widehat{z}+\Lambda+s\phi_{2})^{-(p+1)}\big]\,ds.$
Thus,
$\left|E\left(\phi_{1}\right)-E\left(\phi_{2}\right)\right|+\left|\nabla\left(E\left(\phi_{1}\right)-E\left(\phi_{2}\right)\right)\right|\leq C\left(\left|\phi_{1}-\phi_{2}\right|+\left|\nabla\left(\phi_{1}-\phi_{2}\right)\right|\right).$
For any $t\leq t_{0}$ we have from (69) that
$\big|P(\phi_{1})-P(\phi_{2})\big|+\big|\nabla(P(\phi_{1})-P(\phi_{2}))\big|$
$\qquad\leq C\big(\varepsilon(t)+Be^{\mu t}\big)\Big(|\phi_{1}-\phi_{2}|+|\nabla(\phi_{1}-\phi_{2})|\Big),$
therefore
$\|P(\phi_{1})-P(\phi_{2})\|_{C_{\mu}^{1}((-\infty,t_{0}]\times S^{N-1})}\leq C\big(\varepsilon(t_{0})+Be^{\mu t_{0}}\big)\|\phi_{1}-\phi_{2}\|_{C_{\mu}^{1}((-\infty,t_{0}]\times S^{N-1})}.$
By Theorem 3.7,
$\|\mathcal{T}(\phi_{1})-\mathcal{T}(\phi_{2})\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}$ $\leq C\|P(\phi_{1})-P(\phi_{2})\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})}$
$\leq C\big(\varepsilon(t_{0})+Be^{\mu t_{0}}\big)\|\phi_{1}-\phi_{2}\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}.$
Then we yield (70) by taking $|t_{0}|$ sufficiently large. By the contraction mapping principle, there exists $\phi\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ such that $\mathcal{T}(\phi)=\phi$. This gives a solution $\phi$ to (66). Consequently, $z=\widehat{z}+\phi$ is a solution of $\mathcal{N}(z)=0$ and satisfies (67). ∎
Theorem 4.9.
Let $\eta_{\star}(t)=d_{F}e^{\frac{2p-\alpha}{p+1}t}$ such that $\mathcal{L}_{0}(\eta_{\star})=Fe^{\frac{2p-\alpha}{p+1}t}$.
(1)Radial case. Assume $F\neq 0$ and $\mu>\frac{(2p-\alpha)j}{p+1}$ for some integer $j\geq 2$, and that $\mu\notin\left\{\frac{(2p-\alpha)k}{p+1}:k\geq 2\right\}$. Then there exists $t_{0}<0$ with $|t_{0}|$ sufficiently large, such that there exists a smooth function $\widehat{\varphi}$ on $(-\infty,t_{0}]$ satisfying
$\widehat{z}=\eta_{\star}+\widehat{\varphi}(t),$
and $\widehat{z}$ satisfies the hypotheses of Theorem 4.8.
(2)Nonradial case. Assume that the corresponding index sets $\mathcal{I}_{\rho}$ and $\mathcal{I}_{\widetilde{\rho}}$ given in (37) and (38) (with multiple number ignored) in the proof of Theorem 2.2 respectively. $\mu\notin\mathcal{I}_{\rho}\cup\mathcal{I}_{\widetilde{\rho}}~\text{and}~\mu>\rho_{j+4}>\frac{2p-\alpha}{p+1}$ in cases (31)(32)(34)(35); $\mu>\rho_{1}$in cases (29)(30)(33)(36). Let $\varphi$ be a solution of the equation $\mathcal{L}(\varphi)=0$ on $\mathbb{R}\times\mathbb{S}^{N-1}$, and suppose that $\varphi(t,\theta)\to 0~~\text{as}~t\to-\infty~\text{uniformly}~\text{for}~\theta\in{S^{N-1}}$. Then there exists $t_{0}<0$ with $|t_{0}|$ sufficiently large, such that there exists a smooth function $\widetilde{\varphi}$ on $(-\infty,t_{0}]\times\mathbb{S}^{N-1}$ satisfying
$\widehat{z}=\eta_{\ast}+\varphi+\widetilde{\varphi}$,
and $\widehat{z}$ satisfied the hypotheses of Theorem 4.8.
Proof. We prove case (2); the proof of case (1) is similar and eaiser.
Take a function $\phi(t,\theta)$ such that $\phi(t,\theta)\to 0$ uniformly for $\theta$ on $\mathbb{S}^{N-1}$ as $t\to$-$\infty$. A direct computation yields
$\mathcal{N}(\eta_{\ast}+\phi)=\mathcal{L}(\phi)-\Big[(\eta_{\ast}+\phi+\Lambda)^{-p}-\Lambda^{-p}+p\Lambda^{-(p+1)}(\eta_{\ast}+\phi)\Big]=\mathcal{L}(\phi)-\sum_{i=2}^{\infty}b_{i}(\eta_{\ast}+\phi)^{i}.$ (71)
(For convenience we write full infinite-series form of the nonlinear term ).
We now write the argument for case (31) in the proof of Theorem 2.2. The associated index sets $\mathcal{I}_{\rho}$ and $\mathcal{I}_{\widetilde{\rho}}$ are those given in (37) and (38) (ignoring multiple numbers). In this case,
$\rho_{1}=\sigma_{1}^{(1)},\quad\rho_{2}=\sigma_{1}^{(2)},\ \ldots,\ \rho_{j}=\sigma_{1}^{(j)},\quad\rho_{j+1}=\frac{2p}{p+1},\quad\rho_{j+2}=\sigma_{1}^{(j+1)},\ldots.$
Let $K\geq j+4$ be the largest integer such that $\rho_{K}<\mu$, and let $\widetilde{K}$ be the largest integer such that $\widetilde{\rho}_{\widetilde{K}}<\mu$. Since the kernel of $\mathcal{L}_{0}$ contains no function to zero as $t\to$-$\infty$ .And the kernel of $\mathcal{L}_{k}$ contains one exponentially decaying function $\psi_{k}^{+}(t)=e^{\sigma_{1}^{(k)}t}$ and one exponentially growing function $\psi_{k}^{-}(t)=e^{\sigma_{2}^{(k)}t}$ as . Without loss of generality,we assume that the solution $\varphi$ of $\mathcal{L}(\varphi)=0$ has the form
$\varphi(t,\theta)=\sum_{i=1}^{j}c_{i}Q_{i}(\theta)e^{\rho_{i}t}+\sum_{i=j+2}^{K}c_{i}Q_{\,i-1}(\theta)e^{\rho_{i}t},$ (72)
where $c_{i}$ are constants. This is because any term of the form $e^{\rho_{i}t}$ with $i>K$ appearing in $\varphi$ would contribute only terms $e^{\widetilde{\rho}_{\ell}t}$ with $\widetilde{\rho}_{\ell}>\mu$ in $H(\eta_{\ast}+\varphi)$.
We first consider the case that
$\mathcal{I}_{\rho}\cap\mathcal{I}_{\tilde{\rho}}$$\emptyset$ (73)
We shall prove that it can successively construct $\widetilde{\varphi}_{0},\widetilde{\varphi}_{1},\ldots,\widetilde{\varphi}_{K}$ such that for every $l=0,1,\ldots,\widetilde{K}$,
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=O\big(e^{\widetilde{\rho}_{l+1}t}\big).$ (74)
We first take $\phi=\varphi$, with $\varphi$ given by (72). Then by (71) and the fact that $\mathcal{L}(\varphi)=0$, we have
$\mathcal{N}(\eta_{\ast}+\varphi)=\sum_{n_{1}+\cdots+n_{r_{1}}\geq 2}a_{n_{1}\ldots n_{r_{1}}}e^{(n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}})t}Q_{1}^{n_{1}}\cdots Q_{j}^{n_{j}}Q_{0}^{n_{j+1}}Q_{j+1}^{n_{j+2}}\cdots Q_{r_{1}-1}^{n_{r_{1}}},$
where $n_{1},\ldots,n_{r_{1}}$ are nonnegative integers and $a_{n_{1}\ldots n_{r_{1}}}$ are constants. By the definition of $\mathcal{I}_{\widetilde{\rho}}$ that each $n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}}$ is equal to some $\widetilde{\rho}_{i}$. Hence
$\mathcal{N}(\eta_{\ast}+\varphi)=\sum_{i=1}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{K+1}t}\big),$ (75)
where $\widetilde{M}_{i}$ is given by (41) and $a_{im}$ are constants. In particular,
$\mathcal{N}\left(\eta_{*}+\varphi\right)=O\left(e^{\tilde{\rho}_{1}t}\right),$
Here $\widetilde{\rho}_{1}=2\rho_{1}$. Hence (74) holds when $l=0$ and $\widetilde{\varphi}_{0}=0$.
Assume that we have already constructed $\widetilde{\varphi}_{0},\widetilde{\varphi}_{1},\ldots,\widetilde{\varphi}_{l-1}$ such that (74) holds for $0,1,\ldots,l-1$. We now consider the case $l$. Set
$\widetilde{\varphi}_{l}(t,\theta)=\left(\sum_{m=0}^{\widetilde{M}l}c_{lm}Q_{m}(\theta)\right)e^{\widetilde{\rho}_{l}t},$ (76)
where $c_{lm}$ are constants to be determined. A computation similar to that leading to (75) yields
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=\mathcal{L}(\widetilde{\varphi}_{1})+\cdots+\mathcal{L}(\widetilde{\varphi}_{l})+\sum_{i=1}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big),$
where $a_{im}$ are constants whose values may differ from those in (158). By the induction hypothesis,
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l-1}\big)=\sum_{i=l}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big),$
and hence
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=\mathcal{L}(\widetilde{\varphi}_{l})+\sum_{i=l}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big).$
Note that $\widetilde{\varphi}_{l}$ does not contribute to the coefficients $a_{lm}$. We choose $\widetilde{\varphi}_{l}$ so that
$\mathcal{L}\left(\tilde{\varphi}_{l}\right)=-\left\{\sum_{m=0}^{\tilde{M}_{l}}a_{lm}Q_{m}(\theta)\right\}e^{\tilde{\rho}_{l}t}.$
In this way we obtain (74) for the index $l$. With $\widetilde{\varphi}_{l}$ given by (76), it remains only to solve
$\mathcal{L}_{m}\big(c_{lm}e^{\widetilde{\rho}_{l}t}\big)=-a_{lm}e^{\widetilde{\rho}_{l}t}$ (77)
for $m=0,1,\ldots,\widetilde{M}_{l}$.
Since $\rho_{m}\neq\widetilde{\rho}_{l}$ for all $m,l$, we can find constants $c_{lm}$ satisfying (77). Indeed, we can write an explicit formula for $c_{lm}e^{\widetilde{\rho}_{l}t}$ in terms of $a_{lm}e^{\widetilde{\rho}_{l}t}$ by using the basis of $\operatorname{Ker}(\mathcal{L}_{m})$ given in Lemma 4.3.
if $0<\rho_{m}<\widetilde{\rho}_{l}$, this representation is provided by (63); if $\rho_{m}>\widetilde{\rho}_{l}$, we simply replace the first integral in (63) with the one from $t$ to $t_{0}$. This completes the induction.
In conclusion, we set
$\widetilde{\varphi}(t,\theta)=\sum_{i=1}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}c_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t},$ (78)
where $c_{im}$ are constants. Then
$\mathcal{N}(\eta_{\ast}+\varphi+\widetilde{\varphi})=O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big)=O(e^{\mu t}).$
[…]
We now consider the general situation, in which some $\rho_{i}$ can be written as a positive–integer linear combination of $\rho_{1},\ldots,\rho_{i-1}$. We briefly indicate how to modify the above argument to handle this case. The modification mainly concerns (77). If for some coefficient $a_{lm}$ we have $\rho_{m}=\widetilde{\rho}_{l}$, then instead of choosing only a constant $c_{lm}$ as in (77), we can find constants $c_{l0m}$ and $c_{l1m}$ such that
$\mathcal{L}_{m}\left(\left(c_{l0m}+tc_{l1m}\right)e^{\tilde{\rho}_{l}t}\right)=-a_{lm}e^{\tilde{\rho}_{l}t}.$
Such powers of $t$ will generate higher-order powers of $t$ in the process of the iteration. In general, for a nonnegative integer $J$, if constants $a_{ljm}$ are given for $j=0,1,\ldots,J$, then we can find constants $c_{ljm}$, $j=0,1,\ldots,J+1$, such that
$\mathcal{L}_{m}\left(\sum_{j=0}^{J+1}c_{ljm}t^{j}e^{\widetilde{\rho}_{l}t}\right)=\sum_{j=0}^{J}a_{ljm}t^{j}e^{\widetilde{\rho}_{l}t}.$
Therefore, in the general case we do not adopt (78), but instead take
$\widetilde{\varphi}(t,\theta)=\sum_{i=1}^{\widetilde{K}}\sum_{j=0}^{i}\left\{\sum_{m=0}^{\widetilde{M}_{i}}c_{ijm}Q_{m}(\theta)\right\}t^{j}e^{\widetilde{\rho}_{i}t},$
where $c_{ijm}$ are constants. This completes the proof of the proposition.
In particular, we have now found a function $z(t,\theta)$ satisfying the required conditions. Recalling that
$u(x)=u_{s}(|x|)+|x|^{\frac{\alpha+2}{p+1}}z(\ln|x|,\theta),$
we obtain the desired solution $u$, and the proof of Theorem 1.1 and Thereom 1.2 is thereby completed.
Acknowledgments
Conflict of interest